### v1

# Human2Robot: Learning Robot Actions from Paired Human-Robot Videos

Abstract
Distilling knowledge from human demonstrations is a promising way for robots to learn and act. Existing work often overlooks the differences between humans and robots, producing unsatisfactory results. In this paper, we study how perfectly aligned human-robot pairs benefit robot learning. Capitalizing on VR-based teleportation, we introduce H&R, a third-person dataset with 2,600 episodes, each of which captures the fine-grained correspondence between human hands and robot gripper. Inspired by the recent success of diffusion models, we introduce Human2Robot, an end-to-end diffusion framework that formulates learning from human demonstrates as a generative task. Human2Robot fully explores temporal dynamics in human videos to generate robot videos and predict actions at the same time. Through comprehensive evaluations of 8 seen, changed and unseen tasks in real-world settings, we demonstrate that Human2Robot can not only generate high-quality robot videos but also excel in seen tasks and generalize to unseen objects, backgrounds and even new tasks effortlessly.
1 Introduction
Figure 1: Human2Robot: An end-to-end generative framework that simultaneously predicts robotic videos and actions.
The ability to observe others, whether people or animals, and learn skills to solve new tasks is a key reason humans can handle a wide range of complex challenges. To develop robots capable of assisting with various everyday problems, they must also acquire this skill, particularly the ability to learn directly from humans. This motivates a plethora of work studying how to effectively learn from human demonstrations [5, 58, 39, 59, 4, 55]. However, due to the discrepancies between humans and robots, existing work [59, 39, 70] struggles to transfer learned representations from humans to robots, or simply enforcing consistent outputs between humans and robots [4] without considering the motion trajectories.
In this paper, we argue that fine-grained alignment between humans and robots is of great significance. This is because coarse-grained data provides limited clues for key points and cannot capture detailed information from humans, such as different movement trajectories or grasping positions. However, annotating the detailed correspondence between humans and robots, while feasible [29, 35], is quite time-consuming and labor-intensive. Fortunately, virtual reality (VR)-based teleoperation [28, 15, 45, 13] offers a promising way for collecting synchronized actions of humans and robots thanks to its efficiency and affordability. This motivates us to build upon VR-based teleoperation to study how paired data can help robots learn better. In particular, we focus on obtaining perfectly aligned data between human hands and robot arms as a testbed for learning from human demonstrations. With this in mind, we first introduce a data collection pipeline for gathering paired videos of human hands and robotic arms, aligned frame by frame. Compared to previous work that solely collects demonstration data, our method does not require much additional setup, equipment, or time, making it exceptionally easy to implement. It allows for efficient and consistent data collection. Furthermore, this approach can theoretically be extended to collect data from any type of robotic arms and a wide variety of scenes. Finally, we introduce H&R, a third-person dataset characterized by perfectly aligned videos of human hands and robotic arms, which comprises six fundamental tasks and six long-horizon tasks, totalling 2,600 episodes. Our goal is to train a robot arm that is able to mimic the actions performed by human hands and then perform tasks correspondingly, conditioned on a human demonstration. This requires the robot to know how its arm is aligned with human hands, which is readily available in our dataset, but more importantly, has the ability to act correspondingly. Therefore, the robot needs to predict what actions to take in the future conditioned on human motions as well as historical clues. This is similar in spirit to video generation tasks, which are able to produce perceptually-convincing future frames (actions) based on on a variety of clues [14, 23, 24, 25, 38, 42]. As a result, we build upon video generation models to mimic human actions.
In light of this, we introduce Human2Robot, an end-to-end generative framework that predicts robotic videos and actions to take simultaneously. In particular, Human2Robot is built upon a pretrained Stable Diffusion model, including a spatial UNet for feature extraction, two behavior extractors for motion and position encoding, a spatial-temporal UNet that leverages temporal clues for frame and action prediction. As a result, Human2Robot is not only able to generate high-quality robotic arm videos, which is crucial for developing reliable and explainable models, as it enables users to visually verify and understand the underlying logic of robotic action trajectories, but can also predict corresponding action trajectory required to perform the task. With carefully designed training strategies, Human2Robot not only performs well on routine tasks but also generalizes to different positions, unseen objects, backgrounds and new tasks without additional training. Furthermore, we also introduce a KNN-based approach, enabling Human2Robot to perform routine tasks even without human demonstrations. In summary, our main contributions include:
  • We introduce a low-cost, low-requirement, high-efficiency, and highly scalable pipeline for collecting perfectly aligned data of humans and robots in each frame, and we present H&R, the first dataset featuring perfectly aligned videos of humans and robots.
  • We introduce Human2Robot, a diffusion-based model capable of generating robotic videos from human videos while simultaneously predicting the corresponding robotic trajectories in an end-to-end manner.
  • We propose a KNN+Human2Robot method, which leverages KNN to perform tasks without explicitly requiring human hand videos as a condition.
  • Human2Robot can generate high-quality robotic arm videos, which are crucial for developing reliable and explainable models, as well as excelling in routine tasks, even with variations in position, environment, and objects. Notably, Human2Robot can complete new tasks without any additional training and is capable of executing all tasks in real time.
2 Related Work
Teleoperation. Since current learning-based methods require large amounts of demonstration data, researchers have developed various forms of teleoperation approaches for data collection [34, 21, 54, 45, 57, 2, 3, 47, 19, 71, 17, 1]. However, these approaches may entail specific device requirements and may only be applicable to specific forms of robotic systems. Recently, VR-based methodologies [28, 15] have attracted considerable attention due to their cost-effectiveness, efficiency and versatility. However, these methods map human action information to robot operations and record the results, often overlooking the importance of the paired visual information. We leverage this perception to capture perfectly aligned videos of humans and robots, which are crucial for robotic imitation.
Learning from Human Videos. Owing to the prohibitive expenses of collecting data on real robots, researchers have recently been attempting to leverage existing human-centric video datasets to enhance robot policy learning [35, 56, 10, 53, 70]. Researchers propose to learn representations from human videos to assist robots in task execution [64, 59, 37]. However, these approaches need substantial prior knowledge and struggle to transfer to robots [39, 5]. A recent approach [4] trains robots to make changes to the environment as humans do by simply enforcing consistent outputs between humans and robots. This could overlook crucial details provided by human hands including proper posture and feasible trajectory. Meanwhile, certain methods  [33, 29] have focused on aligning representations across videos, but they do not directly tackle the domain difference between human-centric videos and robot-centric videos. In contrast, our approach utilizes paired data and diffusion models to achieve precise alignment of human and robot videos.
Diffusion Models for Video Generation. Due to the remarkable performance of diffusion models in image generation [48, 49, 41, 51, 6], many approaches have explored their potential for video generation [26, 16, 8, 63, 32, 66]. Moreover, conditionally controlled video generation [40, 30, 60, 27, 65] has broader and more practical application scenarios and has already become a hot topic in academic research. In recent work, several approaches [9, 62, 7] have utilized generative diffusion models to aid policy learning for robots. However, these methods primarily generate videos from a conventional human perspective, often overlooking the distributional differences between human and robotic actions. In contrast, we introduce a novel approach by applying video diffusion models to robotic scenarios, utilizing a generative model conditioned on human behaviors to guide robotic action outputs.
Diffusion Models for Visual Understanding. The diffusion model not only performs impressively in generative domains but also shows promising applications in visual understanding tasks [69, 52, 36, 18]. Some notable approaches involve using diffusion models for object detection [11], image segmentation [67], and visual representation learning [69, 12]. The approach most similar to ours, GenRec [61] proposes a unified video diffusion model that jointly optimizes video generation and video classification, exhibiting the effectiveness of the diffusion model in video understanding. Our approach is designed for robotic tasks, constructing a human-to-robot video generative model that can simultaneously predict actions for subsequent robotic manipulation tasks.
3 H&R Dataset
3.1 Data Collection Pipeline
Our system setup is straightforward, consisting of a Meta Quest 3 VR headset, a 7-DoF xArm robotic arm, and two Intel Realsense D435 cameras capturing images of the human hand and robotic arm from the same viewpoint. The dual-camera setup ensures capturing the synchronized movements of both the human and robotic arms. With the integration of the hardware components, our system aims to achieve a critical objective:
Producing a paired dataset where each data point reflects a seamless synchronization of human hand action and robotic arm response, captured synchronously from identical viewpoints.
Towards this goal, we build our H&R system focusing on developing software and establishing data collection protocols. We build the control system with OpenTeach [28]. When the user puts on the headset and performs several hand operations, the VR headset sends the coordinates of the hand movements of the user to the robotic arm and achieves synchronized actions. Here, one of the biggest challenges lies in the inconsistency of the scale and the orientation between the two coordinate systems. We find a simple yet effective solution to this problem, and the detailed method is provided in the Section B.
3.2 Statistics of H&R Dataset
With our designed system and data collection pipeline, we propose our H&R dataset, the first dataset featuring paired human and robotic videos. Human events vary from six fundamental tasks and six long-horizon tasks. The whole dataset includes 2,600 episodes, and each episode contains frames ranging from 300 to 600. Details can be found in the Section C.1.
To the best of our knowledge, the H&R dataset is the first video dataset that ensures perfectly aligned video between the hand and the robotic arm.
Figure 2: Data demonstration. The image above shows a sample from our dataset, recorded simultaneously by two cameras—one for the human and one for the robot—while the user wears a VR headset to perform the operation using our pipeline.
4 Human2Robot
Figure 3: Architecture overview of Human2Robot. Our approach consists of two training stages. In the first stage, only Behavior Extractors and Spatial Layer in both Spatial UNet and Spatial-Temporal UNet are trainable. In the second stage, we only train Temporal Layer and Policy Head. The Pretrained Encoder and Decoder are frozen all the time. In inference, we have two different modes: the KNN-Based mode can complete seen tasks without hand video, while the Manual Guided mode can complete both seen and unseen tasks. Proprio refers to the proprioception of the robot.
Our goal is to develop a model capable of learning from human demonstrations to predict future robot observations and corresponding actions. Formally, we aim to establish a universal policy $\pi$ that predicts future robot observations $\hat{o}_{r}^{t:t+n}$ and action trajectories $a_{t:t+n}$, leveraging inputs from a sequence of hand demonstration videos $o_{h}^{t:t+n}$, environmental observations $o_{r}^{t}$, and the current proprioception $s_{t}$:
$a_{t:t+n},\ \hat{o}_{r}^{t:t+n}=\pi(o_{h}^{t:t+n},o_{r}^{t},s_{t}),$ (1)
where $n$ represents the length of the action trajectory we aim to predict. This setting requires the model to not only predict the executed actions but also visually simulate future interactions between the robot arm and objects.
4.1 Architecture
Inspired by the recent success of video diffusion models that are able to predict high-fidelity future frames [20], Human2Robot formulates the behavior learning process from human demonstrations as a generative task—generating robot videos and their corresponding actions simultaneously, conditioned on human videos and historical context. Inspired by  [27], Human2Robot explores a Spatial UNet and a Spatial-Temporal UNet, working collaboratively to learn from humans. In particular, the Spatial UNet extracts features from the robot arm, which is further fed into the Spatial-Temporal UNet for temporal modeling. Human2Robot also contains two behavior extractors that estimate position and motion clues using human videos. To achieve action prediction, we use the intermediate features from the Spatial-Temporal UNet to predict the action trajectory via a Policy Head. Figure 3 gives an overview of our pipeline. Below, we introduce the architecture in detail. Spatial UNet. The Spatial UNet, consisting of 4 upsampling blocks and 4 downsampling blocks, extracts useful clues from the robot arm. In particular, each block, known as the Spatial Layer, leverages self-attention to learn features tailored for robot arms, as well as cross-attention with pre-extracted CLIP embedding to incorporate semantic clues. The weights of the Spatial-UNet are initialized from those of the Stable Diffusion [50] to ease the learning process. The features derived by Spatial-UNet provide a condition for predicting future frames as well as actions.
While Spatial UNet introduces a similar number of parameters as the denoising UNet, it only extracts features once during the entire process, unlike diffusion-based video generation where each frame undergoes denoising multiple times. Therefore, it does not significantly increase computational overhead during inference. Behavior Extractor. Recall that we aim to learn from human demonstrations, and thus it is important to extract motion and position information from human videos. This is achieved by Behavior Extractor, which contains four convolution layers (4$\times$ kernels, with a stride of 2). To focus more on the motion trajectories, we employ two independent Behavior Extractors. Behavior Extractor 1 takes the image of the hand corresponding to the robot at the first time step, and Behavior Extractor 2 takes the future image or the entire video of the hand as inputs depending on the training stage, as will be introduced in Section 4.2.
Spatial-Temporal UNet. The Spatial-Temporal UNet aims to predict future frames as well as actions to take simultaneously, requiring the model to learn temporal dynamics. Thus, each block of the Spatial-Temporal UNet (ST-UNet in short) has a Spatial Layer which is the same as the one in Spatial UNet (S-UNet in short), followed by an additional Temporal Layer. More specifically, the ST-UNet takes the features from the behavior extractors and noise latent as inputs. Each Spatial Layer in ST-UNet also concatenates the features of the corresponding layer of the S-UNet, exploring the robot arm as references for prediction. The Temporal Layer focuses on temporal modeling so as to produce high-fidelity future frames as well as accurate action trajectories. Policy Head. We aim to generate future frames as well as actions to take within a single unified framework. Drawing from the insights of [61], we propose to divide the UNet mapping function into two distinct components: $F=F_{tail}\cdot F_{head}$. We treat $F_{head}$ as a feature extractor for motion information. This is achieved by passing the intermediate features from UNet and proprioception data $s_{t}$ to a light-weighted Multilayer Perceptron to predict the actions of the robotic arm. Pretrained Encoder and Decoder. Human2Robot builds upon Stable Diffusion [50], which consists of an encoder and a decoder. The encoder turns images into latent embeddings for fast denoising, and the decoder maps the latents back to images. During training, both the encoder and decoder are kept frozen.
4.2 Training
Our approach consists of two training stages. In the first stage, we focus on image generation by taking the first frame of $o_{r}^{t}$, the first frame of human $o_{h}^{t}$ and future frame of human $o_{h}^{t+i}$ as inputs to predict the future frame of robot $\hat{o}_{r}^{t+i}$. In this stage, we train the S-UNet, the Behavior Extractors, and the ST-UNet without the temporal layers. The spatial layers in both the ST-UNet and S-UNet are instantiated using the powerful open-source Stable Diffusion model (SD) [50], while the Behavior Extractors are initialized with Gaussian weights, except for the final projection layer, which utilizes zero initialized convolution.
In the second stage, we concentrate the training effort on the temporal layer. In this stage, we take a 30-frame segment of human video $o_{h}^{t:t+n}$ and the first frame $o_{r}^{t}$ of the robot video $o_{r}^{t:t+n}$ as inputs to predict the future robot observation $\hat{o}_{r}^{t:t+n}$ and action $a_{t:t+n}$. The first frame of the robot video is fed into Behavior Extractor 1, while the entire video segment is passed into Behavior Extractor 2, yielding features $x_{1}^{t}$ and $x_{2}^{t:t+n}$, respectively. Then, we replicate $x_{1}^{t}$ by n times and add it to $x_{2}^{t:t+n}$, which is subsequently added to noise latent and fed to our ST-UNet. The weights of the VAE Encoder and Decoder, as well as the CLIP image encoder are frozen all the time. During our training process, we find that dense sampling (30fps) results in jittering in generated videos. Therefore, we reduce the frame rate to one-fourth of the original, namely cutting the number of frames by three-quarters. Optimization Objectives. We train Human2Robot with both generation and action prediction objectives, aiming for the model to achieve high-quality video generation and accurate action prediction simultaneously. The goal of generation for the model is to predict the applied noise, so the objective of the optimization process is:
$L_{G}=\mathop{\mathbb{E}_{z_{ts},c,\epsilon,ts}}(||\epsilon-\epsilon_{\theta}(z_{ts},c,ts)||_{2}^{2}),$ (2)
the parameters of the formula represent the latent state $z_{ts}$ at time step $ts$, which is obtained from Spatial-Temporal UNet. The conditioning variable $c$ includes the first frame $o_{r}^{t}$ of the robot and a human video $o_{h}^{t:t+n}$. $\epsilon$ denotes the added noise, and $\epsilon_{\theta}(z_{ts},c,ts)$ represents the model’s predicted noise.
Our action is divided into three components: $a_{t}^{pos}\in\mathbb{R}^{3}$ for 3D position, $a_{t}^{rot}\in\mathbb{R}^{6}$ for rotation and $a_{t}^{grip}\in\{0,1\}$ for gripper state. We use 6D rotation representation proposed in [72, 31] to guarantee continue representation. We use MSE loss $L_{p}$ and $L_{r}$ for position and rotation respectively, and use BCE loss $L_{g}$ for gripper state.
To balance the generation and action prediction losses, we used a set of balancing weights to control the contribution of each loss to the overall objective.
$L=L_{G}+\lambda_{p}L_{p}+\lambda_{r}L_{r}+\lambda_{g}L_{g},$ (3)
where, $\lambda_{p}$, $\lambda_{r}$ and $\lambda_{g}$ are the weights of $L_{p}$, $L_{r}$ and $L_{g}$ respectively.
Method Seen Task Changed Task Unseen Task
Task-1 Task-2 Task-3 Task-4 Task-5 Task-6 Task-7 Task-8
Demonstration Videos Retrieved by KNN
ACT* (KNN) [71] 75 55 0 80 30 - - -
Human2Robot (KNN) 60 85 80 70 60 - - -
Manual Demonstration Videos
ACT* [71] 70 0 0 85 20 45 0 0
Human2Robot 85 95 90 80 75 75 95 85
Table 1: Quantitative evaluation results of KNN and manual guidance. KNN-based methods are not able to complete different tasks with the same initial scene (Task-6) and unseen tasks (Task-7 Task-8). The task description is as follows: Seen Tasks: (1) Pinch cube: pinch the cube and place it on the plate. (2) Push box: push the box from left to right. (3) Push plate: push the plate from top to bottom. Changed Tasks: (4) Pinch a pen on the plate: pinch a pen in a plate and place it on the plate left to it (unseen background, with no plate below the pen in the training set). (5) Pinch a ping pong ball: pinch the ping pong ball (unseen object) and place it on the plate. (6) Pinch a cube to the designated plate (bottom/top): pinch a cube and place it on the bottom/top plate (different trajectories). Unseen Tasks: (7) Pinch a banana from bottom to top: with a plate in the scene (similar to the tasks in the training set), pinch the banana (unseen object) from bottom to top in the image (unseen trajectory). (8) Rotate once before pushing the box: Turn the hand 360° counterclockwise (unseen trajectory) and push the box.
4.3 Inference
With the aforementioned designs, Human2Robot is able to generate high-quality robot videos and perform actions, using observation captured by the camera of the robot and the provided hand video. We further introduce a set of strategies that facilitate more effective task execution by the robot arm, as will be introduced below.
KNN + Human2Robot. To avoid the need to explicitly provide human demonstration videos for every prediction, we use a k-nearest neighbors (KNN) approach to identify the most probable task for the current scene and retrieve the human demonstration video corresponding to the closest matching features as the conditioning input to guide the task execution. Specifically, we use DINOv2 [44] and CLIP [46] as feature extractors to capture features from the first frame of each robotic arm video in the training set. During prediction, we select the $k$ closest features based on the current environment, and the episode with the most frequent and closest match is chosen as the conditioning input, which is depicted in the KNN-Based Inference section of Figure 3.
Using KNN significantly simplifies the process for known tasks. However, for tasks involving different objectives within the same setting—such as Task 6 in Table 1, as well as for unseen tasks, explicit human video guidance is required to accomplish these tasks effectively. Temporal Addition of Actions. To mitigate trajectory jittering during sequential prediction, inspired by the approach used in ACT [71], we derive the action not from a single prediction but from a weighted average of multiple prediction steps. In particular, we predict a series of actions $a_{t:t+n}$ at time t, which are then combined with the previously predicted actions to compute the aggregated actions $A_{t:t+n}^{t}$. This represents the set of actions predicted at time $t$, to be executed from time $t$ to $t+n$, and can be computed as follows:
$A_{t:t-m+n}^{t}=(1-\gamma)A_{t:t-m+n}^{t-m}+\gamma a_{t:t-m+n},$ (4)
$A_{t-m+n:t+n}^{t}=a_{t-m+n:t+n},$ (5)
where $\gamma$ is a hyperparameter, with a larger $\gamma$ indicating a stronger emphasis on the current prediction, and $m$ ($m\leq n$) refers to the length of the action executed after each prediction. Specifically, when $m=n$, the inference degenerates into sequential prediction. Inference Acceleration. During inference, the diffusion model typically goes through $T$ denoising steps to generate the output. We explore the effectiveness of using features from different denoising steps in Section 5.3, and find that features obtained from the first half of the denoising stages lead to better results. Therefore, to balance speed and quality, we halt the denoising process after obtaining features from the $\frac{T}{4}$ step, directly predict and execute the action, without waiting for the entire denoising process to complete. It is worth noting that if we aim to achieve the fastest possible speed, we can directly use the one-step denoising features for prediction, which allows us to predict the next 30 steps of actions within 0.6 seconds.
5 Experiments
5.1 Experimental Setups
Task Definition. We conduct testing across eight experiments, as depicted in Table 1. These experiments are categorized based on their familiarity and complexity into three types: Seen Tasks are present during the training phase but feature slight variations, such as changes in object positions during testing. Changed Tasks are designed to test the generalization capabilities of the model, as it requires the model to understand human hand movements and transfer actions to different types of objects. Unseen Tasks introduce completely new trajectories. Previous methods [71, 39] struggle to achieve this goal because completing a task with an unknown trajectory is highly challenging.
Baseline Construction. We compare Human2Robot with the following baselines. All baselines and Human2Robot utilize the same H&R dataset for training across all tasks for a fair comparison.
  • ACT* [71]: For a fair comparison, we modify ACT to a video conditioned policy, and we name it ACT*. Details can be found in the Section D.
  • XSkill [68]: XSkill is a human-video-conditioned policy managed to bridge the embodiment gap between human and robot.
pinch cube pull plate pinch and pull (unseen)
XSkill [68] 40 60 0
Human2Robot 85 90 60
Table 2: Quantitative results comparison between XSkill and Human2Robot. We compare the basic ability on seen tasks and the generalization ability on unseen tasks (combination of seen tasks).
Evaluation Protocols. For the evaluation of the KNN-based methods, we evaluate models using seen tasks by randomly positioning objects within the defined area during testing. This evaluates the KNN’s effectiveness in identifying and successfully completing the known tasks. For testing scenarios in which the hand demonstration videos are manually given, we collect 10 new data samples, ensuring variation in position, trajectory, and lighting conditions. Additionally, we collect 20 demonstrations for each result to calculate the success rate.
Implementation Details. To demonstrate the effectiveness of our approach, we train our model on H&R. We used 4 NVIDIA A100 GPUs for training in the first stage and 8 4090 GPUs for the second stage. In the first training stage, we use the native image size of 424×240 and conduct training for 15,000 steps and set the total batch size as 128. In the second training stage, the temporal layer and policy header are trained for 50,000 steps using 30-frame video sequences and the total batch size is 8. The learning rates of the two stages are set to 1e-5, and the weights of each objective are set to $\lambda_{p}=\lambda_{r}=\lambda_{g}=1$. During inference, we utilize a DDIM sampler for 30 denoising steps and extract features from the first upsample layer and the 7th denoising step for action prediction. The smoothing parameter $\gamma$ is set to 0.5.
Figure 4: Qualitative visualization of generated result and robot action result.
5.2 Main Results
We discuss the capabilities of the model in generating realistic robot videos and robot manipulation experiments. The evaluation tasks vary from seen to changed and unseen tasks, allowing for a comprehensive study of our method’s imitative ability and generalization.
Action Prediction with KNN. We test the model in a setting where no paired demonstrations are provided. To gain the proper guidance signals, we employ KNN, which utilizes features extracted from DINOv2 [44] and CLIP [46] for searching the high-quality and relevant episodes as hand demonstration videos. As shown in Table 1, Human2Robot performs well across multiple seen tasks, achieving 85% and 80% on tasks 2 and 3, respectively. In comparison, the ACT* model fails to complete these tasks since ACT is not designed for multi-task learning. In particular, due to the prevalence of grasping tasks in our dataset, ACT performs well on grasping tasks (Task-1, Task-4), but poorly on pushing tasks (Task-2, Task-3).
Besides, ACT places more emphasis on the initial scene information than human videos, consistent with its design purpose, which results in it being able to execute tasks based on the initial scene even when the KNN selects an incorrect human hand video. In some cases (Task-1, Task-2, Task-5), its success rate even exceeds that of the ground truth. However, Our model is more reliant on the human hand video, which means that when the KNN selects a less suitable demonstration, the model is likely to fail, leading to poor performance on Task-1.
Moreover, in the Changed Task setting, Human2Robot performs a comparable or better successful rate than ACT* model, achieving 70% on Task-4 and 60% on Task-5, while ACT tends to fail due to the unseen initial frame. Such competitive results demonstrate the high efficiency and accuracy of Human2Robot in action prediction with episodes retrieved by KNN.
Action Prediction with Human Demonstration Videos. We also evaluate performance in scenarios where paired human demonstration videos are provided. As shown in Table 2, Human2Robot outperforms XSkill [68] in both seen and unseen tasks, highlighting the superiority of our model. Besides, as shown in Table 1, Human2Robot significantly benefits from using these GT demonstrations instead of relying solely on KNN, with an improvement in success rates of over 10%. This enhancement is largely attributed to the test set containing some positions not seen during training, which poses a challenge for KNN in finding appropriate demonstrations. Additionally, Human2Robot consistently outperforms ACT* across all seen tasks, unseen tasks, and most changed tasks. Notably, in the Unseen Task category which includes challenging actions such as moving a banana from the bottom to the top of the frame and rotating in a circle before pushing a box—Human2Robot achieves impressive success rates of 95% and 85%, respectively, whereas ACT* struggles to generalize in such condition. These results highlight the strong generalization capabilities of our model, showcasing its ability to perform tasks with significantly greater accuracy than ACT*.
Moreover, on more complex 6DOF task with significant background changes, Human2Robot can also achieve an accuracy of 80%, as shown in Figure 5. For the handwriting task, having one-shot generalization ability for unseen tasks is crucial, as it is impossible to cover all possible characters in the training data. However, to the best of our knowledge, no existing model has achieved this level of generalization—except for Human2Robot, as shown in Figure 6.
Realistic Robot Video Generation. Besides predicting actions, Human2Robot can visually simulate future dynamics of robot movements, requiring precise predictions of the motions of the robot arm and its interactions with objects. As depicted in Figure 4, Human2Robot effectively generates realistic robot arm videos based on the initial state and human demonstrations. Impressively, these videos not only mirror human hand movements but are also consistently accurate in representing real-world robotic movements and interactions. These results highlight the strong imaginative capability of our generative model, demonstrating its ability to bridge visual discrepancies between human and robot arms. It also accurately maps the relationship between human hand movements, robot actions, and object interactions.
Figure 5: Qualitative visualization of upright brush task. A full 6-DOF task with significant background changes and Human2Robot achieves an 80% success rate on this task.
5.3 Discussion
Figure 6: Handwriting task. Trained on random movements data and one-shot to unseen character writing in brand new background.
Generated Video Usage and Evaluation. We further explore the capabilities of Human2Robot by generating demonstrations for a robot-video-conditioned policy using human demonstration videos. In this case, Human2Robot serves to expand the input sources, allowing the use of human videos instead of robot videos, which is more convenient. Specifically, we choose ACT* as the test video-conditioned policy. We use the robot-video-conditioned model, instead of the human-video-conditioned model in Table 1. However, during inference, condition it with robot demonstrations (w Robot GT) and robot demonstrations generated by Human2Robot (w Generated) from human video. As shown in Table 3, the performance of our generated video is competitive with that of the robot video.
Moreover, to further evaluate the quality of our generated video, we treat the above method as a way to assess video quality, measuring the difference between generated and real videos. We regard the performance of robot video as an upper bound. The performance when using generated videos during inference serves as a proxy for evaluating visual quality. As shown in the Table 3 below, replacing GT videos with our generated videos achieves a similar score, demonstrating the realism of our video prediction.
Task1 Task2 Task3 Task4 Task6 Task8 (unseen)
ACT* (w Robot GT) 80 90 0 80 60 0
ACT* (w Generated) 60 80 0 80 40 0
Human2Robot 85 95 90 80 75 85
Table 3: Performance comparison between robot demonstration and robot demonstration generated by Human2Robot from human demonstration.
The Choice of Upsampling Layers for Feature Extraction. As shown in Table 4, we examine the impact of extracting features from different UNet upsample layers for predicting actions via the MLP. Using the second up-sampling block (Up Index 2) achieves a 90% success rate on seen tasks, while the first up-sampling block (Up Index 1) reaches a 95% success rate on unseen tasks. The results indicate that the choice of upsampling layer has minimal effect on outcomes.
Up Index Seen Task Unseen Task
1 85 95
2 90 80
3 80 70
Table 4: Results of different up-sampling blocks.
Denoising Step Seen Task Unseen Task Times(s)
1 80 90 0.6
7 85 95 3.2
15 80 80 6.5
22 70 80 9.3
Table 5: Results of different denoising steps.
Studying the Choice of Denoising Step. As shown in Tab 5, During inference, we fix the upsampling layer to extract features from different denoising steps. We find that, due to the MLP being trained with random noise, features from just one denoising step already allow the model to perform well, achieving 80% accuracy on seen tasks and 90% on unseen tasks, with an inference time of 0.6 seconds for 30 step trajectory prediction. Features after 7 denoising steps (1/4 of the total steps) yield the highest success rate. However, the results are worse at 15 (1/2 total steps) and 22 (3/4 total steps), possibly due to the following reason: in early denoising stages, as shown in Figure 7, the feature focuses more on position information which enhances the success rate. However, in later denoising stages, the model focuses more on reconstructing detailed visual information, which increases action variance, making the movements jerky of the robot and ultimately reducing the success rate.
Figure 7: Visual result of one-step DDIM.
The Effect of Data Scale. After achieving strong image generation capabilities after the first training stage using the whole dataset, we train the model using different data scales in the second stage to assess its performance across varying amounts of data. As shown in Table 6, we conduct tests across various data scales and discover that even a small amount of data can lead to promising results on previously encountered tasks. For instance, using just 5% of the data(116 episodes) is sufficient to achieve a 70% success rate on seen tasks. The success rate on unseen tasks increases significantly as the amount of data increases.
Data Scale Seen Task Unseen Task
0.05 70 30
0.1 70 40
0.2 80 60
1 85 80
Table 6: Results of different data scales.
6 Conclusion
In this paper, we introduced a data collection pipeline for perfectly paired human-robot data and used it to create the H&R dataset. Then we presented Human2Robot, a novel end-to-end diffusion framework that simultaneously predicted robot video and actions given a hand video. Next, we leveraged a KNN method to perform tasks without explicitly requiring human hand videos as input. Human2Robot performed excellently not only on seen tasks but also on changed and unseen tasks and all the tasks can be executed in real time.
Appendix A System Setup Details
The H&R system is equipped with state-of-the-art technology specifically chosen to facilitate seamless and precise interactions between human operators and robotic mechanisms, including a Meta Quest 3 VR headset, a 7-DoF xArm robotic arm, and two Intel Realsense D435 cameras as shown in Figure 9(a) 9(b).
The Meta Quest 3 VR headset is selected for its advanced gesture recognition capability and affordable cost. This gesture recognition capability is integral to our data collection pipeline, providing an efficient and user-friendly way to enhance the natural interaction between human operators and the robotic system. The Xarm robotic arm, which features seven degrees of freedom, provides the flexibility and precision required for complex and varied tasks. This arm is integral to replicating human hand movements accurately, ensuring that the physical actions translated through the VR system are executed precisely by the robotic counterpart. Visual data collection is handled by two Intel Realsense D435 cameras strategically positioned to monitor both human and robot actions. These cameras operate at a resolution of $240\times 424$ pixels, capturing images at 30Hz. The dual-camera setup captures the synchronized movements of both the human and robotic arms. With the integration of the hardware components, our system aims to achieve a critical objective:
Producing a paired dataset where each data point reflects a seamless synchronization of human hand action and robotic arm response, captured synchronously from identical viewpoints.
Toward this goal, we build our H&R system to develop software and establish data collection protocols. We build the control system with OpenTeach [28], connecting different hardware components to enable the operator to perform manual tasks and simultaneously monitor the robotic arm’s actions through the VR headset display. Specifically, we set up two similar environments for human operation and robotic arm operation respectively, as shown in Figure 9, and one camera is placed in each operating space. When the user wears the VR headset and performs hand movements, the headset transmits the coordinates of these movements to the backend, where they are retargeted to control the robotic arm, enabling synchronized actions. Section B illustrates more details about the retargeting process. Here, one of the biggest challenges lies in the inconsistency of the scale and the orientation between the two coordinate systems. In the following, we mainly illustrate our implementation for achieving the coordinate alignment.
Appendix B Data Collection Details
While collecting data, the operating spaces are shown in Figure 9. During the preparation phase, we will record the positional information of four points within the robot’s operating space: three anchor points and the initial position of the robotic arm. The starting position of the robotic arm is recorded and detailed anchor configurations are provided in Section B.1. Operators are required to wear headsets to perform operations throughout the data collection process. The robotic arm will not follow the hand’s movements until the operator sets three anchor points in their workspace by putting hand on each anchor for a brief period. Our program utilizes six anchor points—three for the human and three for the robot—to align the coordinate systems of both sides, as detailed in Section B.1. Since the coordinate system initiated by the headset may vary slightly each time, it is necessary to make a few adjustments to the positions of objects in the environment before starting the actual data collection. During the data collection process, we will use the right hand for demonstration, while the left hand will control the start and end of recording by pressing the keyboard.
B.1 Coordinate System Alignment
In this section, we provide a detailed explanation of how we align the position and rotation of the coordinate systems for the VR headset and the robotic arm.
Figure 8: VR headset Coordinate System.
(a) Human Operating Space.
(b) Robot Operating Space.
(c) Data demonstration.
Figure 9: The operating spaces of human and robot
Position alignment. We utilized the introduced anchor points to align the coordinate systems. As illustrated in Figure 8, we select three anchors within the operating space and derive two mutually perpendicular vectors, $e_{h}^{x}$ and $e_{h}^{y}$ for the human side, and $e_{r}^{x}$ and $e_{r}^{y}$ for the robot side. These are accompanied by the coordinates of the bottom-right anchor point, $o_{h}$ and $o_{r}$, and are subject to the following constraints in the real world, $|e_{h}^{x}|_{w}=|e_{r}^{x}|_{w}$, $|e_{h}^{y}|_{w}=|e_{r}^{y}|_{w}$ and $e_{h}^{x}\cdot e_{h}^{y}=e_{r}^{x}\cdot e_{r}^{y}=0$, where $|\cdot|_{w}$ indicates the real-world distance. From these vectors, we calculate the unit vectors $e_{h}^{z}$ and $e_{r}^{z}$, which are perpendicular to $e_{h}^{x}$ and $e_{h}^{y}$ respectively.
Then, given any hand coordinate $p_{h}$, the corresponding robotic arm coordinate $p_{r}$ can be computed as:
$p_{r}=o_{r}+\mu_{x}e_{r}^{x}+\mu_{y}e_{r}^{y}+\eta\mu_{z}e_{r}^{z},$ (6)
where $\eta$ is a hyperparameter that represents the scaling factor of motion amplitude along the z-axis. The value of $\mu$ can be computed as follows:
$\mu_{i}=\frac{e_{h}^{i}\cdot(p_{h}-o_{h})}{|e_{h}^{i}|^{2}},i\in\{x,y,z\}.$ (7)
Rotation alignment. After aligning the positions, additional alignment is necessary for the orientations. Our goal is to ensure that the relative rotations of the human hand and the robotic arm, relative to their initial states, remain consistent.
To synchronize the rotation of the robotic arm with the human hand, we establish a 3D coordinate system similar to the approach taken by OpenTeach[28]. This coordinate system enables us to compute the rotation matrix $M_{h}^{t}$ for the human hand at time $t$. Additionally, we derive the rotation matrix $M_{r}^{t}$ for the robotic arm at the same time $t$, using the robot’s state parameters such as pitch, yaw, and roll. The rotation matrix $M_{h}^{t}$ and $M_{r}^{t}$ at the time $t$ can be expressed in terms of the rotation matrices $M_{h}^{0}$ and $M_{r}^{0}$ at time $t=0$ as:
$M_{h}^{t}=M_{h}^{0}R_{h}^{t},$ (8)
$M_{r}^{t}=M_{r}^{0}R_{r}^{t},$ (9)
where $R_{h}^{t}$ and $R_{h}^{t}$ represent the rotation relative to the coordinate system $M_{h}^{0}$ and $M_{r}^{0}$ respectively. Recall that our goal is to ensure the relative rotations are the same, which allows us to derive the following equation:
$R_{r}^{t}=P^{-1}R_{h}^{t}P,$ (10)
where $P$ is the transformation matrix defined from the basis of the coordinate system $M_{r}^{0}$ to that of $M_{h}^{0}$.
Organize Equation 8,  9, and 10, and the target rotation matrix of the robotic arm can be calculated as follows:
$M_{r}^{i}=M_{r}^{0}P^{-1}(M_{h}^{0})^{-1}M_{h}^{i}P,$ (11)
(a) Different positions. We positioned the block at two distinct locations in the scene (top-left and top-right of the image) and observed that the model was able to complete the tasks successfully in both cases. Additionally, we placed the box at two different positions on the table, and the model was able to complete the tasks effortlessly.
(b) Unseen object and trajectory. We first tested the model with unseen objects, and the object in the human demonstration video differed from the actual object (top-left image). The results showed that the model successfully completed the task without any noticeable impact. Furthermore, we evaluated the model on grasping tasks involving unseen objects (top-right and bottom-left images). The results demonstrated that the model could successfully complete tasks with objects it had not encountered before. Lastly, we tested the model on unseen tasks with unseen objects (bottom-right image). The results revealed that our model exhibited strong generalization capabilities and successfully completed the task.
Figure 10: Qualitative visualization of generated results and robot actions
Figure 11: Overview of the dataset for base tasks.
Appendix C H&R Details
C.1 Task Details
Our H&R dataset contains 6 base tasks and 6 long-horizon tasks. The definition of tasks in dataset is listed below.
Base tasks. The visual representations of the tasks are shown in Figure 11.
  • Pinch a cup: The cup can initially be placed anywhere on the table. The goal of this task is to pick up the cup and place it in another location.
  • Pinch a cube: The cube comes in two colors (red and green), and it can initially be placed within a designated area. The goal of this task is to pick up the cube and place it on a plate.
  • Pinch a pencil: The pencil comes in two colors (red and black), the goal is to pick up the pencil and place it on a plate.
  • Push a box: The box comes in two colors (red and green), and its initial position of it is random. The goal is to push the box from left to right
  • Pull a plate: The goal is to pull the plate from bottom to top.
  • Push a plate: The goal is to push the plate from top to bottom.
Long-Horizon tasks. These tasks are a combination of the base tasks.
  • Pinch both cubes: There are two red cubes on the table, and the goal is to pick up each cube individually and place them onto the plate.
  • Pinch the designated cube: There is a red cube and a green cube on the table, and the goal is to pick up one of the cubes onto the plate.
  • Pinch to the designated plate: There is a red cube and two plates on the table, and the goal is to pick up the cubes onto one of the plates.
  • Pinch the cube and pull the plate: The combination of pinch a cube and pull the plate. The goal is to first pick the cube onto the plate and pull the plate.
  • Pull the plate and Pinch the cube: The combination of pull the plate and pinch a cube. After pulling the plate, the cube should be placed in the plate’s new position.
  • Return to the initial position: Complete the base tasks and return to the initial position.
C.2 Episode Details
H&R consists of 2,600 episodes, each containing the following information, adhering to the RT-X [43] standards:
  • a paired human hand and robotic arm video containing frames between 200 and 600, with variations depending on the task.
  • state of the robotic arm at each timestamp, including position and rotation.
  • joint of the robotic arm velocity at each timestamp.
  • transform matrix of the human hand, including position and rotation.
  • position of each key point of the human hand.
  • action of the robotic arm retargeted based on human hand data.
  • timestamp.
Appendix D Baseline Details
Here we talk about how we add conditions to ACT [71] in detail. We input the observation information of the robotic arm into the encoder. Then, we add the features of the next $n$ frames of the human hand as a condition to the position embedding and input them into the decoder. Specifically, similar to the method used for processing the original robot images, we input human images into ResNet18 [22] and obtain a feature map of size 300x512. We then pass it through a shared MLP to obtain a feature of size 512, which is added to the position embedding and fed into the Transformer decoder. Both the ResNet18 [22] and MLP are trainable.

### v2

@@
 In this paper, we study how perfectly aligned human-robot pairs benefit robot learning.
-Capitalizing on VR-based teleportation, we introduce H&R, a third-person dataset with 2,600 episodes, each of which captures the fine-grained correspondence between human hands and robot gripper.
-Inspired by the recent success of diffusion models, we introduce Human2Robot, an end-to-end diffusion framework that formulates learning from human demonstrates as a generative task.
+Capitalizing on VR-based teleportation, we introduce H&R, a third-person dataset with 2,600 episodes, each of which captures the fine-grained correspondence between human hand and robot gripper.
+Inspired by the recent success of diffusion models, we introduce Human2Robot, an end-to-end diffusion framework that formulates learning from human demonstration as a generative task.
 Human2Robot fully explores temporal dynamics in human videos to generate robot videos and predict actions at the same time.
-Through comprehensive evaluations of 8 seen, changed and unseen tasks in real-world settings, we demonstrate that Human2Robot can not only generate high-quality robot videos but also excel in seen tasks and generalize to unseen objects, backgrounds and even new tasks effortlessly.
+Through comprehensive evaluations of 4 carefully selected tasks in real-world settings, we demonstrate that Human2Robot can not only generate high-quality robot videos but also excels in seen tasks and generalizing to different positions, unseen appearances, novel instances, and even new backgrounds and task types.
 1 Introduction
@@
 This motivates a plethora of work studying how to effectively learn from human demonstrations [5, 56, 38, 60, 4, 53].
-However, due to the discrepancies between humans and robots, existing work [59, 39, 70] struggles to transfer learned representations from humans to robots, or simply enforcing consistent outputs between humans and robots [4] without considering the motion trajectories.
-In this paper, we argue that fine-grained alignment between humans and robots is of great significance.
-This is because coarse-grained data provides limited clues for key points and cannot capture detailed information from humans, such as different movement trajectories or grasping positions.
-However, annotating the detailed correspondence between humans and robots, while feasible [29, 35], is quite time-consuming and labor-intensive.
-Fortunately, virtual reality (VR)-based teleoperation [28, 15, 45, 13] offers a promising way for collecting synchronized actions of humans and robots thanks to its efficiency and affordability.
-This motivates us to build upon VR-based teleoperation to study how paired data can help robots learn better.
-In particular, we focus on obtaining perfectly aligned data between human hands and robot arms as a testbed for learning from human demonstrations.
-With this in mind, we first introduce a data collection pipeline for gathering paired videos of human hands and robotic arms, aligned frame by frame.
-Compared to previous work that solely collects demonstration data, our method does not require much additional setup, equipment, or time, making it exceptionally easy to implement.
-It allows for efficient and consistent data collection.
-Furthermore, this approach can theoretically be extended to collect data from any type of robotic arms and a wide variety of scenes.
-Finally, we introduce H&R, a third-person dataset characterized by perfectly aligned videos of human hands and robotic arms, which comprises six fundamental tasks and six long-horizon tasks, totalling 2,600 episodes.
-Our goal is to train a robot arm that is able to mimic the actions performed by human hands and then perform tasks correspondingly, conditioned on a human demonstration.
-This requires the robot to know how its arm is aligned with human hands, which is readily available in our dataset, but more importantly, has the ability to act correspondingly.
+However, due to the discrepancies between humans and robots, existing work [60, 38, 71] struggles to transfer learned representations from humans to robots, or simply enforce consistent outputs between humans and robots [4] without considering motion trajectories.
+In this paper, we advocate that fine-grained alignment between humans and robots is of great significance.
+Coarse-grained data provides limited clues for key points and cannot capture detailed information, such as different movement trajectories or grasping positions.
+While annotating detailed correspondences between humans and robots is feasible [29, 34], it is time-consuming and labor-intensive.
+Fortunately, virtual reality(VR)-based teleoperation [28, 15, 43, 13] offers a promising way to collect synchronized actions of humans and robots, making it efficient and affordable.
+This motivates us to build upon VR-based teleoperation for studying paired data that can improve robot learning.
+In particular, we focus on obtaining perfectly aligned data between humans and robots as a testbed for learning from human demonstrations.
+To this end, we introduce a novel data collection pipeline for gathering paired videos of human hands and robotic arms, aligned frame by frame.
+Unlike previous work that solely collects demonstration data, our method requires minimal additional setup, equipment, or time, making it exceptionally easy to implement and facilitating efficient and consistent data collection.
+We alse introduce the H&R dataset, a third-person dataset featuring perfectly aligned videos of human hands and robotic arms, which includes 8 fundamental tasks and 6 long-horizon tasks, totalling 2,600 episodes.
+Our goal is to train a model that mimics human actions and performs corresponding tasks based on human demonstrations.
+This requires the robot to understand how its end effector is aligned with human hands, which is readily available in our dataset, but more importantly, has the ability to act correspondingly.
 Therefore, the robot needs to predict what actions to take in the future conditioned on human motions as well as historical clues.
-This is similar in spirit to video generation tasks, which are able to produce perceptually-convincing future frames (actions) based on on a variety of clues [14, 23, 24, 25, 38, 42].
+This is similar in spirit to video generation tasks, that predict future frames (actions) based on various cues [14, 23, 24, 25, 37, 41].
 As a result, we build upon video generation models to mimic human actions.
@@
 In particular, Human2Robot is built upon a pretrained Stable Diffusion model, including a spatial UNet for feature extraction, two behavior extractors for motion and position encoding, a spatial-temporal UNet that leverages temporal clues for frame and action prediction.
-As a result, Human2Robot is not only able to generate high-quality robotic arm videos, which is crucial for developing reliable and explainable models, as it enables users to visually verify and understand the underlying logic of robotic action trajectories, but can also predict corresponding action trajectory required to perform the task.
-With carefully designed training strategies, Human2Robot not only performs well on routine tasks but also generalizes to different positions, unseen objects, backgrounds and new tasks without additional training.
-Furthermore, we also introduce a KNN-based approach, enabling Human2Robot to perform routine tasks even without human demonstrations.
+As a result, Human2Robot is not only able to generate high-quality robotic videos, which facilitating the expansion of robot-conditioned policies to human-conditioned ones, but can also predict corresponding action trajectory required to perform the task.
+With carefully designed training strategies, Human2Robot not only excels on seen tasks but also generalizes to different positions, unseen appearances, instances and even new backgrounds and task types without additional training.
+Furthermore, we introduce a KNN-based approach, enabling Human2Robot to perform seen tasks even without human demonstrations, providing a flexible and scalable solution to robotic learning from demonstrations.
 In summary, our main contributions include:
-  • We introduce a low-cost, low-requirement, high-efficiency, and highly scalable pipeline for collecting perfectly aligned data of humans and robots in each frame, and we present H&R, the first dataset featuring perfectly aligned videos of humans and robots.
-  • We introduce Human2Robot, a diffusion-based model capable of generating robotic videos from human videos while simultaneously predicting the corresponding robotic trajectories in an end-to-end manner.
-  • We propose a KNN+Human2Robot method, which leverages KNN to perform tasks without explicitly requiring human hand videos as a condition.
-  • Human2Robot can generate high-quality robotic arm videos, which are crucial for developing reliable and explainable models, as well as excelling in routine tasks, even with variations in position, environment, and objects.
-Notably, Human2Robot can complete new tasks without any additional training and is capable of executing all tasks in real time.
+  • We introduce a novel low-cost, low-requirement, high-efficiency, and highly scalable pipeline for collecting perfectly aligned data of humans and robots at the frame level.
+Additionally, we present H&R, the first dataset featuring perfectly aligned videos of human hands and robotic arms across a variety of tasks, enabling high-fidelity learning from human demonstrations.
+  • We introduce Human2Robot, an end-to-end generative framework built upon a diffusion-based model, capable of generating robotic videos from human videos while simultaneously predicting the corresponding robotic action trajectories.
+This approach combines state-of-the-art video generation techniques with multi-task action prediction, facilitating accurate task execution.
+  • We propose a KNN+Human2Robot method, which integrates KNN for task prediction, enabling to perform tasks even without human hand videos as input.
+This further enhances the scalability and flexibility of the system.
+  • Human2Robot can generate high-quality robotic arm videos, which facilitates the expansion of robot-conditioned policy to human-conditioned policy, while also excelling in four carefully selected tasks, even with variations in positions, appearances, instances, backgrounds and different task types.
 2 Related Work
 Teleoperation.
-Since current learning-based methods require large amounts of demonstration data, researchers have developed various forms of teleoperation approaches for data collection [34, 21, 54, 45, 57, 2, 3, 47, 19, 71, 17, 1].
-However, these approaches may entail specific device requirements and may only be applicable to specific forms of robotic systems.
+Researchers have developed various forms of teleoperation approaches for data collection [33, 21, 52, 43, 55, 2, 3, 45, 19, 72, 17, 1].
+However, these approaches may entail specific device requirements.
 Recently, VR-based methodologies [28, 15] have attracted considerable attention due to their cost-effectiveness, efficiency and versatility.
-However, these methods map human action information to robot operations and record the results, often overlooking the importance of the paired visual information.
+However, these methods often overlook the importance of the paired visual information.
 We leverage this perception to capture perfectly aligned videos of humans and robots, which are crucial for robotic imitation.
 Learning from Human Videos.
-Owing to the prohibitive expenses of collecting data on real robots, researchers have recently been attempting to leverage existing human-centric video datasets to enhance robot policy learning [35, 56, 10, 53, 70].
-Researchers propose to learn representations from human videos to assist robots in task execution [64, 59, 37].
-However, these approaches need substantial prior knowledge and struggle to transfer to robots [39, 5].
-A recent approach [4] trains robots to make changes to the environment as humans do by simply enforcing consistent outputs between humans and robots.
-This could overlook crucial details provided by human hands including proper posture and feasible trajectory.
-Meanwhile, certain methods  [33, 29] have focused on aligning representations across videos, but they do not directly tackle the domain difference between human-centric videos and robot-centric videos.
-In contrast, our approach utilizes paired data and diffusion models to achieve precise alignment of human and robot videos.
+Researchers have recently been attempting to leverage existing human-centric video datasets to enhance robot policy learning [34, 54, 10, 51, 71].
+Researchers propose to learn representations from human videos to assist in task execution [65, 60, 36].
+However, these approaches need substantial prior knowledge and struggle to transfer to robots [38, 5, 4].
+Meanwhile, some human-video-conditioned methods  [7, 29, 69] have focused on aligning representations across human and robot videos, they are neither efficient nor capable of generalization.
+In contrast, our approach utilizes paired data and diffusion models to achieve strong generalization capabilities.
 Diffusion Models for Video Generation.
-Due to the remarkable performance of diffusion models in image generation [48, 49, 41, 51, 6], many approaches have explored their potential for video generation [26, 16, 8, 63, 32, 66].
-Moreover, conditionally controlled video generation [40, 30, 60, 27, 65] has broader and more practical application scenarios and has already become a hot topic in academic research.
+Due to the remarkable performance of diffusion models in image generation [46, 47, 40, 49, 6], many approaches have explored their potential for video generation [26, 16, 8, 64, 32, 67, 57, 58, 59].
+Moreover, conditionally controlled video generation [39, 30, 61, 27, 66] has broader and more practical application scenarios.
 In recent work, several approaches [9, 63, 7] have utilized generative diffusion models to aid policy learning for robots.
@@
 The approach most similar to ours, GenRec [62] proposes a unified video diffusion model that jointly optimizes video generation and video classification, exhibiting the effectiveness of the diffusion model in video understanding.
-Our approach is designed for robotic tasks, constructing a human-to-robot video generative model that can simultaneously predict actions for subsequent robotic manipulation tasks.
+Our approach is designed for robotic tasks, constructing a human-to-robot video generative model that can simultaneously predict actions for manipulation tasks.
 3 H&R Dataset
 3.1 Data Collection Pipeline
-Our system setup is straightforward, consisting of a Meta Quest 3 VR headset, a 7-DoF xArm robotic arm, and two Intel Realsense D435 cameras capturing images of the human hand and robotic arm from the same viewpoint.
+Our system setup is straightforward, consisting of a Meta Quest 3 VR headset, a 6-DoF xArm robotic arm with a xArm gripper, and two Intel Realsense D435 cameras capturing images of the human hand and robotic arm from the same viewpoint.
 The dual-camera setup ensures capturing the synchronized movements of both the human and robotic arms.
-With the integration of the hardware components, our system aims to achieve a critical objective:
-Producing a paired dataset where each data point reflects a seamless synchronization of human hand action and robotic arm response, captured synchronously from identical viewpoints.
+With the integration of the hardware components, our system aims to achieve a critical objective: Producing a paired dataset where each data point reflects a seamless synchronization of human hand action and robotic arm response, captured synchronously from identical viewpoints.
 Towards this goal, we build our H&R system focusing on developing software and establishing data collection protocols.
@@
 When the user puts on the headset and performs several hand operations, the VR headset sends the coordinates of the hand movements of the user to the robotic arm and achieves synchronized actions.
-Here, one of the biggest challenges lies in the inconsistency of the scale and the orientation between the two coordinate systems.
-We find a simple yet effective solution to this problem, and the detailed method is provided in the Section B.
+Achieving this functionality requires addressing issues such as coordinate alignment, precision, and jitter. The details of the implementation are provided in the Supplementary materials B.2.
 3.2 Statistics of H&R Dataset
 With our designed system and data collection pipeline, we propose our H&R dataset, the first dataset featuring paired human and robotic videos.
-Human events vary from six fundamental tasks and six long-horizon tasks.
+Human events vary from 8 fundamental tasks and 6 long-horizon tasks.
 The whole dataset includes 2,600 episodes, and each episode contains frames ranging from 300 to 600.
-Details can be found in the Section C.1.
-To the best of our knowledge, the H&R dataset is the first video dataset that ensures perfectly aligned video between the hand and the robotic arm.
+Details can be found in Supplementary materials C.
+To the best of our knowledge, the H&R dataset is the first video dataset that ensures perfectly aligned video between human and robot.
 Figure 2: Data demonstration.
@@
 Our goal is to develop a model capable of learning from human demonstrations to predict future robot observations and corresponding actions.
-Formally, we aim to establish a universal policy $\pi$ that predicts future robot observations $\hat{o}_{r}^{t:t+n}$ and action trajectories $a_{t:t+n}$, leveraging inputs from a sequence of hand demonstration videos $o_{h}^{t:t+n}$, environmental observations $o_{r}^{t}$, and the current proprioception $s_{t}$:
-$a_{t:t+n},\ \hat{o}_{r}^{t:t+n}=\pi(o_{h}^{t:t+n},o_{r}^{t},s_{t}),$ (1)
+Formally, we aim to establish a universal policy $\pi$ that predicts future robot observations $o^{r}_{t:t+n}$ and action trajectories $a_{t:t+n}$, leveraging inputs from a sequence of hand demonstration videos $o^{h}_{t:t+n}$, environmental observations $o^{r}_{t}$, and the current proprioception $s_{t}$:
+$a_{t:t+n},\ o^{r}_{t:t+n}=\pi(o^{h}_{t:t+n},o^{r}_{t},s_{t}),$ (1)
 where $n$ represents the length of the action trajectory we aim to predict.
@@
 Our approach consists of two training stages.
-In the first stage, we focus on image generation by taking the first frame of $o_{r}^{t}$, the first frame of human $o_{h}^{t}$ and future frame of human $o_{h}^{t+i}$ as inputs to predict the future frame of robot $\hat{o}_{r}^{t+i}$.
+In the first stage, we focus on image generation by taking the first frame of $o^{r}_{t}$, the first frame of human $o^{h}_{t}$ and future frame of human $o^{h}_{t+i}$ as inputs to predict the future frame of robot $o^{r}_{t+i}$.
 In this stage, we train the S-UNet, the Behavior Extractors, and the ST-UNet without the temporal layers.
@@
 In the second stage, we concentrate the training effort on the temporal layer.
-In this stage, we take a 30-frame segment of human video $o_{h}^{t:t+n}$ and the first frame $o_{r}^{t}$ of the robot video $o_{r}^{t:t+n}$ as inputs to predict the future robot observation $\hat{o}_{r}^{t:t+n}$ and action $a_{t:t+n}$.
+In this stage, we take a 30-frame segment of human video $o^{h}_{t:t+n}$ and the first frame $o^{r}_{t}$ of the robot video $o^{r}_{t:t+n}$ as inputs to predict the future robot observation $o^{r}_{t:t+n}$ and action $a_{t:t+n}$.
 The first frame of the robot video is fed into Behavior Extractor 1, while the entire video segment is passed into Behavior Extractor 2, yielding features $x_{1}^{t}$ and $x_{2}^{t:t+n}$, respectively.
@@
 the parameters of the formula represent the latent state $z_{ts}$ at time step $ts$, which is obtained from Spatial-Temporal UNet.
-The conditioning variable $c$ includes the first frame $o_{r}^{t}$ of the robot and a human video $o_{h}^{t:t+n}$.
+The conditioning variable $c$ includes the first frame $o^{r}_{t}$ of the robot and a human video $o^{h}_{t:t+n}$.
 $\epsilon$ denotes the added noise, and $\epsilon_{\theta}(z_{ts},c,ts)$ represents the model’s predicted noise.
@@
 where, $\lambda_{p}$, $\lambda_{r}$ and $\lambda_{g}$ are the weights of $L_{p}$, $L_{r}$ and $L_{g}$ respectively.
-Method Seen Task Changed Task Unseen Task
-Task-1 Task-2 Task-3 Task-4 Task-5 Task-6 Task-7 Task-8
-Demonstration Videos Retrieved by KNN
-ACT* (KNN) [71] 75 55 0 80 30 - - -
-Human2Robot (KNN) 60 85 80 70 60 - - -
-Manual Demonstration Videos
-ACT* [71] 70 0 0 85 20 45 0 0
-Human2Robot 85 95 90 80 75 75 95 85
-Table 1: Quantitative evaluation results of KNN and manual guidance.
-KNN-based methods are not able to complete different tasks with the same initial scene (Task-6) and unseen tasks (Task-7 Task-8).
-The task description is as follows: Seen Tasks: (1) Pinch cube: pinch the cube and place it on the plate.
-(2) Push box: push the box from left to right.
-(3) Push plate: push the plate from top to bottom.
-Changed Tasks: (4) Pinch a pen on the plate: pinch a pen in a plate and place it on the plate left to it (unseen background, with no plate below the pen in the training set).
-(5) Pinch a ping pong ball: pinch the ping pong ball (unseen object) and place it on the plate.
-(6) Pinch a cube to the designated plate (bottom/top): pinch a cube and place it on the bottom/top plate (different trajectories).
-Unseen Tasks: (7) Pinch a banana from bottom to top: with a plate in the scene (similar to the tasks in the training set), pinch the banana (unseen object) from bottom to top in the image (unseen trajectory).
-(8) Rotate once before pushing the box: Turn the hand 360° counterclockwise (unseen trajectory) and push the box.
 4.3 Inference
@@
 Using KNN significantly simplifies the process for known tasks.
-However, for tasks involving different objectives within the same setting—such as Task 6 in Table 1, as well as for unseen tasks, explicit human video guidance is required to accomplish these tasks effectively.
+However, for tasks involving different objectives within the same setting—such as pick cube task in Figure 4 and Table 2, as well as for unseen tasks such as handwriting, explicit human video guidance is required to accomplish these tasks effectively.
 Temporal Addition of Actions.
@@
 It is worth noting that if we aim to achieve the fastest possible speed, we can directly use the one-step denoising features for prediction, which allows us to predict the next 30 steps of actions within 0.6 seconds.
+Figure 4: Qualitative results. Visualization of generated result and robot action result.
 5 Experiments
 5.1 Experimental Setups
-Task Definition.
-We conduct testing across eight experiments, as depicted in Table 1.
-These experiments are categorized based on their familiarity and complexity into three types: Seen Tasks are present during the training phase but feature slight variations, such as changes in object positions during testing.
-Changed Tasks are designed to test the generalization capabilities of the model, as it requires the model to understand human hand movements and transfer actions to different types of objects.
-Unseen Tasks introduce completely new trajectories.
-Previous methods [71, 39] struggle to achieve this goal because completing a task with an unknown trajectory is highly challenging.
-Baseline Construction. We compare Human2Robot with the following baselines. All baselines and Human2Robot utilize the same H&R dataset for training across all tasks for a fair comparison.
-  • ACT* [71]: For a fair comparison, we modify ACT to a video conditioned policy, and we name it ACT*. Details can be found in the Section D.
-  • XSkill [68]: XSkill is a human-video-conditioned policy managed to bridge the embodiment gap between human and robot.
-pinch cube pull plate pinch and pull (unseen)
-XSkill [68] 40 60 0
-Human2Robot 85 90 60
-Table 2: Quantitative results comparison between XSkill and Human2Robot. We compare the basic ability on seen tasks and the generalization ability on unseen tasks (combination of seen tasks).
-Evaluation Protocols.
-For the evaluation of the KNN-based methods, we evaluate models using seen tasks by randomly positioning objects within the defined area during testing.
-This evaluates the KNN’s effectiveness in identifying and successfully completing the known tasks.
-For testing scenarios in which the hand demonstration videos are manually given, we collect 10 new data samples, ensuring variation in position, trajectory, and lighting conditions.
-Additionally, we collect 20 demonstrations for each result to calculate the success rate.
-Implementation Details.
-To demonstrate the effectiveness of our approach, we train our model on H&R.
-We used 4 NVIDIA A100 GPUs for training in the first stage and 8 4090 GPUs for the second stage.
-In the first training stage, we use the native image size of 424×240 and conduct training for 15,000 steps and set the total batch size as 128.
-In the second training stage, the temporal layer and policy header are trained for 50,000 steps using 30-frame video sequences and the total batch size is 8.
-The learning rates of the two stages are set to 1e-5, and the weights of each objective are set to $\lambda_{p}=\lambda_{r}=\lambda_{g}=1$.
-During inference, we utilize a DDIM sampler for 30 denoising steps and extract features from the first upsample layer and the 7th denoising step for action prediction.
-The smoothing parameter $\gamma$ is set to 0.5.
-Figure 4: Qualitative visualization of generated result and robot action result.
+Task Definition. As visualized in Figure 4, we conduct tests across the following four tasks:
+  • Push the box. Push the box from left to right.
+  • Pick the cube to the designated plate. Pick up the cube and place it on the bottom or top plate depending on the human demonstration.
+  • Pick up and hold the brush upright. Pick up the brush, hold it upright, and get ready to write.
+  • Handwriting. Write the corresponding character depending on the human demonstration.
+The four carefully selected tasks above are designed to test the capabilities in multiple aspects: (1) The push box task is to test the accuracy in predicting height, position, and trajectory.
+(2) The pick cube task requires the model to understand the task objective in the video.
+(3) The upright task is a challenging fully 6-DOF task.
+(4) The handwriting task is to test the generalization ability.
+Notably, in the handwriting task, the models are evaluated on unseen characters.
+Since there are countless characters in the world, it is impossible to demonstrate all of them.
+Therefore, having such a generalization ability is crucial.
+Thus, this task demands exceptional generalization capabilities, which we have not observed in other models.
+Baselines. We compare with the following baselines:
+  • RVCD Policy: We design a robot-video-conditioned policy $\pi(a_{t}|s_{t},v)$, where $v$ refers to the robot video that indicates the task.
+Specifically, we modify ACT [72] to a video conditioned policy.
+Details can be found in Supplementary materials D.2.
+  • RVCD Policy w.
+Generated video: The framework and weights of this model is the same as RVCD Policy, but the robot video, generated by Human2Robot using human demonstration, is provided during inference.
+  • XSkill [69]: XSkill is a human-video-conditioned policy tries to bridge the embodiment gap between human and robot. Details can be found in Supplementary materials D.2.
+  • Human2Robot w. KNN Human2Robot with our KNN method proposed in Section 4.3, which enables to perform tasks without explicit demonstrations.
+The baselines and Human2Robot utilize the same H&R dataset for training across all tasks except handwriting.
+For the handwriting task, the models are trained on handwriting play data [60] (randomly moving data).
+Tasks Push Box Pick and Place Upright Handwriting
+Apperence Position Instance Background Bg and Task
+Generalization Type Brown Blue Pos1 Pos2 Pingpong Banana Bg1 Bg2 H R
+XSkill [69]
+Human2Robot
+Table 1: Generalization evaluation.
+We evaluate generalization across four different tasks in unseen settings, including variations in appearances, positions, instances, backgrounds, and task targets.
+The scenes for each task are shown in the figure above.
+Detailed process can be found in the supplementary materials C.
+Each setting is evaluated with one trial.
 5.2 Main results
-We discuss the capabilities of the model in generating realistic robot videos and robot manipulation experiments.
-The evaluation tasks vary from seen to changed and unseen tasks, allowing for a comprehensive study of our method’s imitative ability and generalization.
-Action Prediction with KNN.
-We test the model in a setting where no paired demonstrations are provided.
-To gain the proper guidance signals, we employ KNN, which utilizes features extracted from DINOv2 [44] and CLIP [46] for searching the high-quality and relevant episodes as hand demonstration videos.
-As shown in Table 1, Human2Robot performs well across multiple seen tasks, achieving 85% and 80% on tasks 2 and 3, respectively.
-In comparison, the ACT* model fails to complete these tasks since ACT is not designed for multi-task learning.
-In particular, due to the prevalence of grasping tasks in our dataset, ACT performs well on grasping tasks (Task-1, Task-4), but poorly on pushing tasks (Task-2, Task-3).
-Besides, ACT places more emphasis on the initial scene information than human videos, consistent with its design purpose, which results in it being able to execute tasks based on the initial scene even when the KNN selects an incorrect human hand video.
-In some cases (Task-1, Task-2, Task-5), its success rate even exceeds that of the ground truth.
-However, Our model is more reliant on the human hand video, which means that when the KNN selects a less suitable demonstration, the model is likely to fail, leading to poor performance on Task-1.
-Moreover, in the Changed Task setting, Human2Robot performs a comparable or better successful rate than ACT* model, achieving 70% on Task-4 and 60% on Task-5, while ACT tends to fail due to the unseen initial frame.
-Such competitive results demonstrate the high efficiency and accuracy of Human2Robot in action prediction with episodes retrieved by KNN.
-Action Prediction with Human Demonstration Videos.
-We also evaluate performance in scenarios where paired human demonstration videos are provided.
-As shown in Table 2, Human2Robot outperforms XSkill [68] in both seen and unseen tasks, highlighting the superiority of our model.
-Besides, as shown in Table 1, Human2Robot significantly benefits from using these GT demonstrations instead of relying solely on KNN, with an improvement in success rates of over 10%.
-This enhancement is largely attributed to the test set containing some positions not seen during training, which poses a challenge for KNN in finding appropriate demonstrations.
-Additionally, Human2Robot consistently outperforms ACT* across all seen tasks, unseen tasks, and most changed tasks.
-Notably, in the Unseen Task category which includes challenging actions such as moving a banana from the bottom to the top of the frame and rotating in a circle before pushing a box—Human2Robot achieves impressive success rates of 95% and 85%, respectively, whereas ACT* struggles to generalize in such condition.
-These results highlight the strong generalization capabilities of our model, showcasing its ability to perform tasks with significantly greater accuracy than ACT*.
-Moreover, on more complex 6DOF task with significant background changes, Human2Robot can also achieve an accuracy of 80%, as shown in Figure 5.
-For the handwriting task, having one-shot generalization ability for unseen tasks is crucial, as it is impossible to cover all possible characters in the training data.
-However, to the best of our knowledge, no existing model has achieved this level of generalization—except for Human2Robot, as shown in Figure 6.
-Realistic Robot Video Generation.
+Push Box Pick Cube Upright Handwriting
+RVCD w. Gen 60 30 35 0
+RVCD 75 45 40 0
+XSkill [69] 70 40 50 0
+Human2Robot w. KNN 85 40 60 0
+Human2Robot 95 85 80 90
+Table 2: Main results. Each task is evaluated with 20 trails.
+Effectiveness.
+As shown in Table 2, Human2Robot achieves a high success rate across all tasks.
+In the pick cube to the designated plate task, half of the test cases require placing the cube on the bottom plate, while the other half require placing it on the top plate.
+Therefore, a success rate above 50% indicates that the model can correctly follow human demonstrations to perform distinct tasks.
+However, all baselines perform poorly on this task, consistently placing the cube on the bottom plate regardless of the given demonstration.
+This proves that only Human2Robot can truly understand human demonstration and execute the correct tasks.
+In addition, all baselines can successfully complete the handwriting task that requires strong generalization exhibitting they have limited generalization capabilities.
+These results highlight the satisfactory performance and strong generalization capabilities of Human2Robot.
+We also test our KNN method proposed in Section 4.3, in a setting where no demonstrations are given.
+The performance of Human2Robot w.
+KNN is shown in Table 2.
+Human2Robot with KNN outperforms all other baselines across all tasks, demonstrating that even without direct demonstrations, Human2Robot can still achieve strong performance.
+However, the KNN method experiences a 10-20% drop in success rate on the push box and upright tasks, which is within an acceptable range.
+Meanwhile, a significant performance drop is observed in the pick cube to the designated plate task.
+This is expected, as the KNN method struggles to select the desired demonstration video when faced with the same initial scene.
+Finally, the KNN method cannot select demonstrations that were not present during training, as these demonstrations are only used during inference.
+As a result, it fails to accomplish the handwriting task.
+As a whole, such competitive results demonstrate the high efficiency and accuracy of Human2Robot in action prediction with episodes retrieved by KNN.
+Quality of Generated Videos.
 Besides predicting actions, Human2Robot can visually simulate future dynamics of robot movements, requiring precise predictions of the motions of the robot arm and its interactions with objects.
-As depicted in Figure 4, Human2Robot effectively generates realistic robot arm videos based on the initial state and human demonstrations.
-Impressively, these videos not only mirror human hand movements but are also consistently accurate in representing real-world robotic movements and interactions.
-These results highlight the strong imaginative capability of our generative model, demonstrating its ability to bridge visual discrepancies between human and robot arms.
-It also accurately maps the relationship between human hand movements, robot actions, and object interactions.
-Figure 5: Qualitative visualization of upright brush task. A full 6-DOF task with significant background changes and Human2Robot achieves an 80% success rate on this task.
+The visual quality of the generated video of Human2Robot is shown in Figure 4.
+However, to evaluate the quality of generated videos, the traditional methods of using video distribution scores presents the following challenges: (1) Since the test set is small, which is common in the robotics domain, typically containing only 5 to 20 test samples, the scores may be unreliable.
+(2) No existing generation methods are directly applicable to our task.
+To this end, we introduce a new visual quality evaluation pipeline: training a RVCD policy as an upper bound and replacing robot-conditioned videos with generated videos during inference.
+The performance drop serves as a proxy for evaluating visual quality.
+As shown in Table 2, the performance of RVCD w.
+Gen achieve a similar score, demonstrating the realism of our video prediction.
+5.3 Generalization
+Position generalization.
+Human2Robot uses its generative capability to enable the policy to understand human intentions and perform corresponding tasks at different positions.
+As shown in Table 1, Human2Robot achieves better results in different testing positions.
+Appearance generalization.
+Appearance generalization is crucial for the model’s robustness.
+Since Human2Robot is based on Stable Diffusion [48], which contains a vast amount of prior knowledge, it naturally inherits strong generalization capabilities in terms of appearance without any data augmentation.
+As shown in the Table 1, Human2Robot demonstrates excellent generalization across different colors, which are unseen during training.
+Instance generalization.
+Instance generalization requires the ability to generalize across objects of different sizes and shapes, making it even more critical than appearance generalization.
+As illustrated in the Table 1, thanks to the generative capability of diffusion models, Human2Robot achieves strong performance across various objects.
+Human2Robot successfully completes tasks on unseen instances, demonstrating its powerful generalization ability.
+Background generalization.
+Although the training data primarily consists of white backgrounds, as shown in the Figure 4, we find that Human2Robot maintains stable performance even when the background color changes or the scene becomes cluttered.
+As presented in the Table 1, Human2Robot successfully completes tasks across different background, demonstrating its strong capability.
+Task generalization.
+We train the model on play data, which involves moving randomly on the table rather than writing any specific character, and then evaluate its one-shot task execution ability given human handwriting prompts of never-before-seen character.
+To keep the scene clean during data collection, no actual handwriting was performed, resulting in blank backgrounds in the training data, which differ from those in the evaluation phase.
+Consequently, the handwriting task simultaneously challenges the generalization ability of models in both scene adaptation and task execution.
+The satisfying results shown in the Table 1 demonstrate the effectiveness and generalization capability of Human2Robot in task transfer.
 5.4 Discussion
-Figure 6: Handwriting task. Trained on random movements data and one-shot to unseen character writing in brand new background.
-Generated Video Usage and Evaluation.
-We further explore the capabilities of Human2Robot by generating demonstrations for a robot-video-conditioned policy using human demonstration videos.
-In this case, Human2Robot serves to expand the input sources, allowing the use of human videos instead of robot videos, which is more convenient.
-Specifically, we choose ACT* as the test video-conditioned policy.
-We use the robot-video-conditioned model, instead of the human-video-conditioned model in Table 1.
-However, during inference, condition it with robot demonstrations (w Robot GT) and robot demonstrations generated by Human2Robot (w Generated) from human video.
-As shown in Table 3, the performance of our generated video is competitive with that of the robot video.
-Moreover, to further evaluate the quality of our generated video, we treat the above method as a way to assess video quality, measuring the difference between generated and real videos.
-We regard the performance of robot video as an upper bound.
-The performance when using generated videos during inference serves as a proxy for evaluating visual quality.
-As shown in the Table 3 below, replacing GT videos with our generated videos achieves a similar score, demonstrating the realism of our video prediction.
-Task1 Task2 Task3 Task4 Task6 Task8 (unseen)
-ACT* (w Robot GT) 80 90 0 80 60 0
-ACT* (w Generated) 60 80 0 80 40 0
-Human2Robot 85 95 90 80 75 85
-Table 3: Performance comparison between robot demonstration and robot demonstration generated by Human2Robot from human demonstration.
 The Choice of Upsampling Layers for Feature Extraction.
-As shown in Table 4, we examine the impact of extracting features from different UNet upsample layers for predicting actions via the MLP.
-Using the second up-sampling block (Up Index 2) achieves a 90% success rate on seen tasks, while the first up-sampling block (Up Index 1) reaches a 95% success rate on unseen tasks.
+As shown in Table 3, we examine the impact of extracting features from different UNet upsample layers for action prediction.
+Using the 2nd up-sampling block (Up Index 2) achieves a 90% success rate on seen tasks, while the 1st up-sampling block (Up Index 1) reaches a 95% success rate on unseen tasks.
 The results indicate that the choice of upsampling layer has minimal effect on outcomes.
@@
 Table 3: Results of different up-sampling blocks.
+Studying the Choice of Denoising Step.
+As shown in Tab 4, During inference, we fix the upsampling layer to extract features from different denoising steps.
+We find that, due to the MLP being trained with random noise, features from just one denoising step already allow the model to perform well, achieving 80% accuracy on seen tasks and 90% on unseen tasks, with an inference time of 0.6 seconds for 30 step trajectory prediction.
+Features after 7 denoising steps (1/4 of the total steps) yield the highest success rate.
+However, the results are worse at 15 (1/2 total steps) and 22 (3/4 total steps), possibly due to the following reason: the features in the early denoising stages, such as the 1st and 7th denoising stage shown in Figure 5, already capture sufficient spatial information to complete the tasks.
+However, in later denoising stages, the model focuses more on reconstructing detailed visual information, which increases action variance, making the movements jerky of the robot and ultimately reducing the success rate.
 Denoising Step Seen Task Unseen Task Times(s)
@@
 Table 4: Results of different denoising steps.
-Studying the Choice of Denoising Step.
-As shown in Tab 5, During inference, we fix the upsampling layer to extract features from different denoising steps.
-We find that, due to the MLP being trained with random noise, features from just one denoising step already allow the model to perform well, achieving 80% accuracy on seen tasks and 90% on unseen tasks, with an inference time of 0.6 seconds for 30 step trajectory prediction.
-Features after 7 denoising steps (1/4 of the total steps) yield the highest success rate.
-However, the results are worse at 15 (1/2 total steps) and 22 (3/4 total steps), possibly due to the following reason: in early denoising stages, as shown in Figure 7, the feature focuses more on position information which enhances the success rate.
-However, in later denoising stages, the model focuses more on reconstructing detailed visual information, which increases action variance, making the movements jerky of the robot and ultimately reducing the success rate.
-Figure 7: Visual result of one-step DDIM.
+Figure 5: Visual result of 1 and 7 step DDIM.
 The Effect of Data Scale.
@@
 Next, we leveraged a KNN method to perform tasks without explicitly requiring human hand videos as input.
-Human2Robot performed excellently not only on seen tasks but also on changed and unseen tasks and all the tasks can be executed in real time.
+Human2Robot not only achieved outstanding performance on four carefully selected tasks but also demonstrated its capabilities in five generalization ability tests.
 Appendix A System Setup Details
-The H&R system is equipped with state-of-the-art technology specifically chosen to facilitate seamless and precise interactions between human operators and robotic mechanisms, including a Meta Quest 3 VR headset, a 7-DoF xArm robotic arm, and two Intel Realsense D435 cameras as shown in Figure 9(a) 9(b).
+The H&R system is equipped with state-of-the-art technology specifically chosen to facilitate seamless and precise interactions between human operators and robotic mechanisms, including a Meta Quest 3 VR headset, a 7-DoF xArm robotic arm, and two Intel Realsense D435 cameras as shown in Figure 7(a) 7(b).
 The Meta Quest 3 VR headset is selected for its advanced gesture recognition capability and affordable cost.
@@
 $M_{r}^{i}=M_{r}^{0}P^{-1}(M_{h}^{0})^{-1}M_{h}^{i}P,$ (11)
-(a) Different positions.
-We positioned the block at two distinct locations in the scene (top-left and top-right of the image) and observed that the model was able to complete the tasks successfully in both cases.
-Additionally, we placed the box at two different positions on the table, and the model was able to complete the tasks effortlessly.
-(b) Unseen object and trajectory.
-We first tested the model with unseen objects, and the object in the human demonstration video differed from the actual object (top-left image).
-The results showed that the model successfully completed the task without any noticeable impact.
-Furthermore, we evaluated the model on grasping tasks involving unseen objects (top-right and bottom-left images).
-The results demonstrated that the model could successfully complete tasks with objects it had not encountered before.
-Lastly, we tested the model on unseen tasks with unseen objects (bottom-right image).
-The results revealed that our model exhibited strong generalization capabilities and successfully completed the task.
-Figure 10: Qualitative visualization of generated results and robot actions
 Figure 8: Overview of the dataset for base tasks.
+B.2 System stability
+Accuracy.
+There are two main factors affecting control accuracy: one is the precision of the data source, i.e., the accuracy of hand gesture recognition, and the other is the accuracy loss caused by mapping hand movements to the robotic arm.
+Our gesture recognition data comes from the Meta Quest 3, and we found that under sufficient lighting conditions, it provides highly accurate hand tracking.
+A traditional mapping approach is to map the wrist position to the end-effector position of robot.
+However, this creates a topological inconsistency, leading to accuracy errors.
+A more topologically consistent approach is to map the midpoint between the thumb and index fingertip.
+While this method preserves structural consistency, it introduces jitter due to the differing speeds of the thumb and index finger when grasping objects.
+Thus, we track the Metacarpophalangeal Joint (MCP) of the index finger, ensuring both the prevention of jitter and the maintenance of accuracy.
+Jitter Handling.
+We set the control mode to serial, making the system’s sensitivity depend on execution speed.
+By allowing 100–300ms latency, we trade speed for stability, reducing jitter and ensuring smoother motion.
+Figure 9: Visualization of generalization evaluation process. To accelerate inference, we skip the video generation process when performing the handwriting task.
 Appendix C H&R Details
 C.1 Task Details
-Our H&R dataset contains 6 base tasks and 6 long-horizon tasks. The definition of tasks in dataset is listed below.
+Our H&R dataset contains 8 base tasks and 6 long-horizon tasks. The definition of our tasks in our dataset is listed below.
 Base tasks. The visual representations of the tasks are shown in Figure 8.
-  • Pinch a cup: The cup can initially be placed anywhere on the table. The goal of this task is to pick up the cup and place it in another location.
-  • Pinch a cube: The cube comes in two colors (red and green), and it can initially be placed within a designated area. The goal of this task is to pick up the cube and place it on a plate.
-  • Pinch a pencil: The pencil comes in two colors (red and black), the goal is to pick up the pencil and place it on a plate.
-  • Push a box: The box comes in two colors (red and green), and its initial position of it is random. The goal is to push the box from left to right
-  • Pull a plate: The goal is to pull the plate from bottom to top.
-  • Push a plate: The goal is to push the plate from top to bottom.
+  • Pick and place the cup: The cup can initially be placed anywhere on the table. The goal of this task is to pick up the cup and place it in another location.
+  • Pick and place the cube: The cube comes in two colors (red and green), and it can initially be placed within a designated area.
+The goal of this task is to pick up the cube and place it on a plate.
+  • Pick and place pencil: The pencil comes in two colors (red and black), the goal is to pick up the pencil and place it on a plate.
+  • Push the box: The box comes in two colors (red and green), and its initial position of it is random. The goal is to push the box from left to right
+  • Pull the plate: The goal is to pull the plate from bottom to top.
+  • Push the plate: The goal is to push the plate from top to bottom.
+  • Pick up and hold the brush upright: The goal is to pick up the brush and upright it.
+  • Handwriting: Play data, write aimlessly on a blank desk.
 Long-Horizon tasks. These tasks are a combination of the base tasks.
-  • Pinch both cubes: There are two red cubes on the table, and the goal is to pick up each cube individually and place them onto the plate.
-  • Pinch the designated cube: There is a red cube and a green cube on the table, and the goal is to pick up one of the cubes onto the plate.
-  • Pinch to the designated plate: There is a red cube and two plates on the table, and the goal is to pick up the cubes onto one of the plates.
-  • Pinch the cube and pull the plate: The combination of pinch a cube and pull the plate. The goal is to first pick the cube onto the plate and pull the plate.
-  • Pull the plate and Pinch the cube: The combination of pull the plate and pinch a cube. After pulling the plate, the cube should be placed in the plate’s new position.
+  • Pick both cubes: There are two red cubes on the table, and the goal is to pick up each cube individually and place them onto the plate.
+  • Pick the designated cube: There is a red cube and a green cube on the table, and the goal is to pick up one of the cubes onto the plate.
+  • Pick to the designated plate: There is a red cube and two plates on the table, and the goal is to pick up the cubes onto one of the plates.
+  • Pick the cube and pull the plate: The combination of pinch a cube and pull the plate. The goal is to first pick the cube onto the plate and pull the plate.
+  • Pull the plate and pick the cube: The combination of pull the plate and pinch a cube. After pulling the plate, the cube should be placed in the plate’s new position.
   • Return to the initial position: Complete the base tasks and return to the initial position.
 C.2 Episode Details
-H&R consists of 2,600 episodes, each containing the following information, adhering to the RT-X [43] standards:
+H&R consists of 2,600 episodes, each containing the following information, adhering to the RT-X standards:
   • a paired human hand and robotic arm video containing frames between 200 and 600, with variations depending on the task.
@@
   • timestamp.
-Appendix D Baseline Details
+Appendix D Experimental Details
+D.1 Experimental Details
+Implementation Details.
+To demonstrate the effectiveness of our approach, we train our model on H&R.
+We used 4 NVIDIA A100 GPUs for training in the first stage and 8 4090 GPUs for the second stage.
+In the first training stage, we use the native image size of 424×240 and conduct training for 15,000 steps and set the total batch size as 128.
+In the second training stage, the temporal layer and policy header are trained for 50,000 steps using 30-frame video sequences and the total batch size is 8.
+The learning rates of the two stages are set to 1e-5, and the weights of each objective are set to $\lambda_{p}=\lambda_{r}=\lambda_{g}=1$.
+During inference, we utilize a DDIM sampler for 30 denoising steps and extract features from the first upsample layer and the 7th denoising step for action prediction.
+The smoothing parameter $\gamma$ is set to 0.5.
+D.2 Baseline Details
+RVCD Policy.
 Here we talk about how we add conditions to ACT [72] in detail.
@@
 Both the ResNet18 [22] and MLP are trainable.
+XSkill.
+We replicated xSkill in a real-world setting, where the authors did not fully open-source the code.
+As a result, we completed the remaining code based on the code from the simulation environment.
+During the skill discovery and labeled dataset phases, we used the default settings, except for adjusting the num frames parameter of the frame sampler from 100 to 50 to reduce GPU memory consumption.
+We train this stage for 300 epochs.
+In the skill transfer composing phase, we found that performance of XSkill was quite poor.
+Although the behavior cloning loss decreased significantly, the diffusion policy barely demonstrated any task execution ability after denoising.
+Since the XSkill [69] paper mentions that any imitation learning policy can be used with the XSkill framework, we replaced the diffusion policy with an action head the same as Human2Robot, a simple MLP, for training.
+During training, we adjusted the prediction horizon to 32, resized the pretraining to 160x120 to maintain consistency with the previous training process, and set the vision feature dimension to 64, in line with the simulation environment.
+We trained this stage for a total of 2000 epochs until the loss showed no significant changes.

### v3

@@
 Distilling knowledge from human demonstrations is a promising way for robots to learn and act.
-Existing work often overlooks the differences between humans and robots, producing unsatisfactory results.
-In this paper, we study how perfectly aligned human-robot pairs benefit robot learning.
-Capitalizing on VR-based teleportation, we introduce H&R, a third-person dataset with 2,600 episodes, each of which captures the fine-grained correspondence between human hand and robot gripper.
-Inspired by the recent success of diffusion models, we introduce Human2Robot, an end-to-end diffusion framework that formulates learning from human demonstration as a generative task.
-Human2Robot fully explores temporal dynamics in human videos to generate robot videos and predict actions at the same time.
-Through comprehensive evaluations of 4 carefully selected tasks in real-world settings, we demonstrate that Human2Robot can not only generate high-quality robot videos but also excels in seen tasks and generalizing to different positions, unseen appearances, novel instances, and even new backgrounds and task types.
+Existing methods, which often rely on coarsely-aligned video pairs, are typically constrained to learning global or task-level features.
+As a result, they tend to neglect the fine-grained frame-level dynamics required for complex manipulation and generalization to novel tasks.
+We posit that this limitation stems from a vicious circle of inadequate datasets and the methods they inspire.
+To break this cycle, we propose a paradigm shift that treats fine-grained human-robot alignment as a conditional video generation problem.
+To this end, we first introduce H&R, a novel third-person dataset containing 2,600 episodes of precisely synchronized human and robot motions, collected using a VR teleoperation system.
+We then present Human2Robot, a framework designed to leverage this data.
+Human2Robot employs a Video Prediction Model to learn a rich and implicit representation of robot dynamics by generating robot videos from human input, which in turn guides a decoupled action decoder.
+Our real-world experiments demonstrate that this approach not only achieves high performance on seen tasks but also exhibits significant one-shot generalization to novel positions, objects, instances, and even new task categories.
 1 Introduction
-Figure 1: Human2Robot: An end-to-end generative framework that simultaneously predicts robotic videos and actions.
-The ability to observe others, whether people or animals, and learn skills to solve new tasks is a key reason humans can handle a wide range of complex challenges.
-To develop robots capable of assisting with various everyday problems, they must also acquire this skill, particularly the ability to learn directly from humans.
-This motivates a plethora of work studying how to effectively learn from human demonstrations [5, 56, 38, 60, 4, 53].
-However, due to the discrepancies between humans and robots, existing work [60, 38, 71] struggles to transfer learned representations from humans to robots, or simply enforce consistent outputs between humans and robots [4] without considering motion trajectories.
-In this paper, we advocate that fine-grained alignment between humans and robots is of great significance.
-Coarse-grained data provides limited clues for key points and cannot capture detailed information, such as different movement trajectories or grasping positions.
-While annotating detailed correspondences between humans and robots is feasible [29, 34], it is time-consuming and labor-intensive.
-Fortunately, virtual reality(VR)-based teleoperation [28, 15, 43, 13] offers a promising way to collect synchronized actions of humans and robots, making it efficient and affordable.
-This motivates us to build upon VR-based teleoperation for studying paired data that can improve robot learning.
-In particular, we focus on obtaining perfectly aligned data between humans and robots as a testbed for learning from human demonstrations.
-To this end, we introduce a novel data collection pipeline for gathering paired videos of human hands and robotic arms, aligned frame by frame.
-Unlike previous work that solely collects demonstration data, our method requires minimal additional setup, equipment, or time, making it exceptionally easy to implement and facilitating efficient and consistent data collection.
-We alse introduce the H&R dataset, a third-person dataset featuring perfectly aligned videos of human hands and robotic arms, which includes 8 fundamental tasks and 6 long-horizon tasks, totalling 2,600 episodes.
-Our goal is to train a model that mimics human actions and performs corresponding tasks based on human demonstrations.
-This requires the robot to understand how its end effector is aligned with human hands, which is readily available in our dataset, but more importantly, has the ability to act correspondingly.
-Therefore, the robot needs to predict what actions to take in the future conditioned on human motions as well as historical clues.
-This is similar in spirit to video generation tasks, that predict future frames (actions) based on various cues [14, 23, 24, 25, 37, 41].
-As a result, we build upon video generation models to mimic human actions.
-In light of this, we introduce Human2Robot, an end-to-end generative framework that predicts robotic videos and actions to take simultaneously.
-In particular, Human2Robot is built upon a pretrained Stable Diffusion model, including a spatial UNet for feature extraction, two behavior extractors for motion and position encoding, a spatial-temporal UNet that leverages temporal clues for frame and action prediction.
-As a result, Human2Robot is not only able to generate high-quality robotic videos, which facilitating the expansion of robot-conditioned policies to human-conditioned ones, but can also predict corresponding action trajectory required to perform the task.
-With carefully designed training strategies, Human2Robot not only excels on seen tasks but also generalizes to different positions, unseen appearances, instances and even new backgrounds and task types without additional training.
-Furthermore, we introduce a KNN-based approach, enabling Human2Robot to perform seen tasks even without human demonstrations, providing a flexible and scalable solution to robotic learning from demonstrations.
+Figure 1: Human2Robot: An human-video-conditioned policy, capable of completing seen tasks and one-shot performing unseen tasks with a single human video.
+The ability to observe others, whether humans or animals, and acquire skills to solve new tasks is a fundamental reason why humans excel at tackling a wide range of complex challenges.
+To enable robots to assist with diverse real-world problems, it is essential that they develop a similar capability, particularly the ability to learn directly from human demonstrations.
+This has motivated extensive research (Bahl et al.
+2023; Srirama et al.
+2024; Nair et al.
+2022; Wang et al.
+2023; Bahl, Gupta, and Pathak 2022; Smith et al.
+2019) on how to learn from human-provided demonstrations effectively.
+Despite recent progress, current approaches face a fundamental generalization gap.
+While they perform well on tasks encountered during training, they often fail entirely when presented with human demonstrations of unseen tasks.
+The prevailing paradigm relies on applying self-supervised (Xu et al.
+2023b) or contrastive learning (Jain et al.
+2024) methods to coarsely aligned human-robot video pairs.
+However, the lack of fine-grained supervision in existing datasets fundamentally constrains what models can learn.
+Rather than uncovering detailed action-level correspondences, models are often limited to capturing global features or high-level task summaries.
+Consequently, many approaches extract holistic video representations using models such as the Perceiver Resampler (Alayrac et al.
+2022), which compresses entire clips into fixed-length embeddings.
+This severely limits their ability to model the nuanced frame-level temporal dynamics that are essential for generalization.
+Furthermore, even when densely aligned data are available, models that rely on global feature matching remain fundamentally limited in their ability to capture fine-grained spatio-temporal structure.
+They often discard the frame-by-frame dynamics that are essential for imitating complex tasks.
+This limitation has created a vicious circle: the lack of fine-grained datasets gives rise to methods that are incapable of leveraging detailed supervision, while the dominance of such methods in turn discourages the development of fine-grained datasets needed to overcome this barrier.
+As such, we argue that to achieve true generalization where a robot can perform tasks demonstrated by a human but unseen in its own training data, we must break this cycle.
+This necessitates a paradigm shift in both learning methodology and data curation.
+In this paper, we propose that the key to unlocking fine-grained human-robot alignment lies in video generation (Ho et al.
+2022; Esser et al.
+2023; Blattmann et al.
+2023a; Wu et al.
+2023; Khachatryan et al.
+2023; Xing et al.
+2024; Tu et al.
+2024a).
+Instead of merely learning whether two videos are broadly similar or mapping task goals from the human domain to the robot domain, we aim to generate a corresponding robot video directly from a human demonstration.
+We posit that by training a model to predict the precise frame-by-frame evolution of a robot’s movements, it learns a far richer and more temporally coherent alignment.
+This generative process forces the model to internalize the intricate dynamics of manipulation, enabling it to understand how a task is done, not just what is done.
+To train such a model would require a dataset with densely aligned and fine-grained human-robot video pairs.
+Manually annotating such data is prohibitively labor-intensive.
+To overcome this, we leverage virtual reality teleoperation to create a novel dataset, which we call H&R.
+By enhancing existing teleoperation systems with improved coordinate system matching, we achieve a seamless mapping between the operator’s hand movements and the robot arm’s motion.
+This allows us to efficiently collect a large-scale third-person dataset of 2,600 episodes, featuring perfectly synchronized videos of human hands and robot arms across 4 types of basic tasks and 6 long-horizon tasks.
+Building on the proposed dataset, we introduce Human2Robot, a framework designed to leverage the proposed H&R dataset.
+It employs a two-stage training process.
+First, we train a Video Prediction Model (VPM) built upon a pretrained Stable Diffusion model, which learns to translate a human video into a robot video.
+This model includes a spatial UNet for feature extraction, a behavior extractors for motion and position encoding, a spatial-temporal UNet that uses a spatial-temporal UNet architecture to explicitly model motion and temporal dynamics, generating a rich latent representation that captures core dynamics of the task.
+Second, we train an action decoder that conditions on the predictive representations generated by the VPM to output robot actions.
+This two-stage design effectively learn human-robot alignment and leverages the implicitly learned robot-dynamics features to guide the final policy learning.
+With this carefully designed framework, Human2Robot not only excels on seen tasks but also demonstrates remarkable one-shot generalization to novel object positions, appearances, and even entirely new task types and backgrounds.
+Furthermore, we introduce a KNN-based inference method that allows Human2Robot to perform previously seen tasks with high precision, even without a human demonstration at test time.
 In summary, our main contributions include:
-  • We introduce a novel low-cost, low-requirement, high-efficiency, and highly scalable pipeline for collecting perfectly aligned data of humans and robots at the frame level.
-Additionally, we present H&R, the first dataset featuring perfectly aligned videos of human hands and robotic arms across a variety of tasks, enabling high-fidelity learning from human demonstrations.
-  • We introduce Human2Robot, an end-to-end generative framework built upon a diffusion-based model, capable of generating robotic videos from human videos while simultaneously predicting the corresponding robotic action trajectories.
-This approach combines state-of-the-art video generation techniques with multi-task action prediction, facilitating accurate task execution.
-  • We propose a KNN+Human2Robot method, which integrates KNN for task prediction, enabling to perform tasks even without human hand videos as input.
+  • We present H&R, the first dataset featuring perfectly aligned videos of human hands and robotic arms across a variety of tasks, enabling high-fidelity learning from human demonstrations.
+  • We introduce Human2Robot, an end-to-end generative framework utilizes a two-stage training process, which excels in carefully selected tasks, even with variations in positions, appearances, instances, backgrounds and different task types.
+  • We propose a KNN+Human2Robot method, which integrates KNN for task prediction, enabling to perform tasks even without human videos as input.
 This further enhances the scalability and flexibility of the system.
-  • Human2Robot can generate high-quality robotic arm videos, which facilitates the expansion of robot-conditioned policy to human-conditioned policy, while also excelling in four carefully selected tasks, even with variations in positions, appearances, instances, backgrounds and different task types.
 2 Related Work
 Teleoperation.
-Researchers have developed various forms of teleoperation approaches for data collection [33, 21, 52, 43, 55, 2, 3, 45, 19, 72, 17, 1].
-However, these approaches may entail specific device requirements.
-Recently, VR-based methodologies [28, 15] have attracted considerable attention due to their cost-effectiveness, efficiency and versatility.
-However, these methods often overlook the importance of the paired visual information.
-We leverage this perception to capture perfectly aligned videos of humans and robots, which are crucial for robotic imitation.
+Recently, VR-based methodologies (Iyer et al.
+2024; Ding et al.
+2024) have attracted considerable attention due to their cost-effectiveness, efficiency and versatility.
+However, these methods focus on controlling the robot, whereas our goal of using VR is to capture perfectly aligned videos of humans and robots, which are essential for robotic imitation.
 Learning from Human Videos.
-Researchers have recently been attempting to leverage existing human-centric video datasets to enhance robot policy learning [34, 54, 10, 51, 71].
-Researchers propose to learn representations from human videos to assist in task execution [65, 60, 36].
-However, these approaches need substantial prior knowledge and struggle to transfer to robots [38, 5, 4].
-Meanwhile, some human-video-conditioned methods  [7, 29, 69] have focused on aligning representations across human and robot videos, they are neither efficient nor capable of generalization.
-In contrast, our approach utilizes paired data and diffusion models to achieve strong generalization capabilities.
+Researchers have recently been attempting to leverage existing human-centric video datasets to enhance robot policy learning (Liu et al.
+2018; Smith et al.
+2020; Chen, Nair, and Finn 2021; Shaw, Bahl, and Pathak 2022; Zeng et al.
+2024).
+Researchers propose to learn representations from human videos to assist in task execution (Xiao et al.
+2022; Wang et al.
+2023; Majumdar et al.
+2023).
+However, these approaches need substantial prior knowledge and struggle to transfer to robots (Nair et al.
+2022; Bahl et al.
+2023; Bahl, Gupta, and Pathak 2022).
+Meanwhile, some human-video-conditioned methods  (Bharadhwaj et al.
+2024; Jain et al.
+2024; Xu et al.
+2023b) have focused on aligning representations across human and robot videos, they are neither efficient nor capable of generalization.
+In contrast, our approach uses paired data and diffusion models to achieve strong generalization capabilities.
 Diffusion Models for Video Generation.
-Due to the remarkable performance of diffusion models in image generation [46, 47, 40, 49, 6], many approaches have explored their potential for video generation [26, 16, 8, 64, 32, 67, 57, 58, 59].
-Moreover, conditionally controlled video generation [39, 30, 61, 27, 66] has broader and more practical application scenarios.
-In recent work, several approaches [9, 63, 7] have utilized generative diffusion models to aid policy learning for robots.
-However, these methods primarily generate videos from a conventional human perspective, often overlooking the distributional differences between human and robotic actions.
-In contrast, we introduce a novel approach by applying video diffusion models to robotic scenarios, utilizing a generative model conditioned on human behaviors to guide robotic action outputs.
-Diffusion Models for Visual Understanding.
-The diffusion model not only performs impressively in generative domains but also shows promising applications in visual understanding tasks [70, 50, 35, 18].
-Some notable approaches involve using diffusion models for object detection [11], image segmentation [68], and visual representation learning [70, 12].
-The approach most similar to ours, GenRec [62] proposes a unified video diffusion model that jointly optimizes video generation and video classification, exhibiting the effectiveness of the diffusion model in video understanding.
-Our approach is designed for robotic tasks, constructing a human-to-robot video generative model that can simultaneously predict actions for manipulation tasks.
+Current research has achieved remarkable performance of video generation (Ho et al.
+2022; Esser et al.
+2023; Blattmann et al.
+2023a; Wu et al.
+2023; Khachatryan et al.
+2023; Xing et al.
+2024; Tu et al.
+2024a, b, c).
+Moreover, diffusion models for video stylization (Ye et al.
+2025; Liu et al.
+2023) and human image animation (Hu 2024; Tu et al.
+2024a) have also shown its promising potential for visual and motion transfer.
+Moreover, diffusion model not only performs impressively in generative domains but also shows promising applications in visual understanding task.
+Some notable approaches involve using diffusion models for object detection (Chen et al.
+2023), image segmentation (Xu et al.
+2023a), and visual representation learning (Yu et al.
+2024; Weng et al.
+2024).
+Thus, we aim to utilize the strong capabilities of diffusion model to learn human-robot alignment and leverages the implicitly learned robot-dynamics features to guide the final policy learning.
 3 H&R Dataset
-3.1 Data Collection Pipeline
-Our system setup is straightforward, consisting of a Meta Quest 3 VR headset, a 6-DoF xArm robotic arm with a xArm gripper, and two Intel Realsense D435 cameras capturing images of the human hand and robotic arm from the same viewpoint.
-The dual-camera setup ensures capturing the synchronized movements of both the human and robotic arms.
-With the integration of the hardware components, our system aims to achieve a critical objective: Producing a paired dataset where each data point reflects a seamless synchronization of human hand action and robotic arm response, captured synchronously from identical viewpoints.
-Towards this goal, we build our H&R system focusing on developing software and establishing data collection protocols.
-We build the control system with OpenTeach [28].
-When the user puts on the headset and performs several hand operations, the VR headset sends the coordinates of the hand movements of the user to the robotic arm and achieves synchronized actions.
-Achieving this functionality requires addressing issues such as coordinate alignment, precision, and jitter. The details of the implementation are provided in the Supplementary materials B.2.
+3.1 Coordinate Alignment for Teleoperation
+Although teleoperation has become a mainstream method for data collection, no one has yet attempted to use it for collecting paired data of humans and robots.
+We aim to provide a feasible solution for this.
+We identify two core issues.
+The first issue is that the coordinate system of the robotic arm is not aligned with that of the human hand, which results in a mismatch between the movements of the human hand and the robotic arm in the video.
+For example, the hand may move across the entire screen, while the robotic arm only moves halfway.
+This visual difference could lead to difficulties in policy learning.
+The second issue is the embodiment gap between the human hand and the robotic gripper, which makes teleoperation unsuitable for certain tasks, such as screwing.
+To address the issue of coordinate system alignment, we record corresponding three points in the real world of the robotic arm and human hand.
+We leverage these points to establish a shared coordinate system with consistent scale for both the robotic arm and the human hand.
+This enables the movement range of the robotic arm to match that of the human hand, as illustrated in Figure 1.
+Details on how to establish a shared coordinate system can be found in the Supplementary Materials.
+However, the embodiment gap remains a challenging issue for current teleoperation system, and we leave it as future work.
+Therefore, at the data level, we focus on collecting pick-and-place data with human and robotic arm alignment.
+In our experiments, we evaluate how well the model generalizes after being trained on relatively simple data.
 3.2 Statistics of H&R Dataset
-With our designed system and data collection pipeline, we propose our H&R dataset, the first dataset featuring paired human and robotic videos.
-Human events vary from 8 fundamental tasks and 6 long-horizon tasks.
+Based on our teleoperation method, we propose our H&R dataset, the first dataset featuring paired human and robot videos.
+Human events vary from 4 types of basic tasks and 6 long-horizon tasks, as shown in Figure 2.
 The whole dataset includes 2,600 episodes, and each episode contains frames ranging from 300 to 600.
-Details can be found in Supplementary materials C.
 To the best of our knowledge, the H&R dataset is the first video dataset that ensures perfectly aligned video between human and robot.
-Figure 2: Data demonstration.
-The image above shows a sample from our dataset, recorded simultaneously by two cameras—one for the human and one for the robot—while the user wears a VR headset to perform the operation using our pipeline.
+Figure 2: Dataset Overview. (L) The ratio of four basic task types and long tasks. (R) Platform environment and the object instances used.
 4 Human2Robot
@@
 Our approach consists of two training stages.
-In the first stage, only Behavior Extractors and Spatial Layer in both Spatial UNet and Spatial-Temporal UNet are trainable.
-In the second stage, we only train Temporal Layer and Policy Head.
-The Pretrained Encoder and Decoder are frozen all the time.
-In inference, we have two different modes: the KNN-Based mode can complete seen tasks without hand video, while the Manual Guided mode can complete both seen and unseen tasks.
-Proprio refers to the proprioception of the robot.
-Our goal is to develop a model capable of learning from human demonstrations to predict future robot observations and corresponding actions.
-Formally, we aim to establish a universal policy $\pi$ that predicts future robot observations $o^{r}_{t:t+n}$ and action trajectories $a_{t:t+n}$, leveraging inputs from a sequence of hand demonstration videos $o^{h}_{t:t+n}$, environmental observations $o^{r}_{t}$, and the current proprioception $s_{t}$:
-$a_{t:t+n},\ o^{r}_{t:t+n}=\pi(o^{h}_{t:t+n},o^{r}_{t},s_{t}),$ (1)
-where $n$ represents the length of the action trajectory we aim to predict.
-This setting requires the model to not only predict the executed actions but also visually simulate future interactions between the robot arm and objects.
-4.1 Architecture
-Inspired by the recent success of video diffusion models that are able to predict high-fidelity future frames [20], Human2Robot formulates the behavior learning process from human demonstrations as a generative task—generating robot videos and their corresponding actions simultaneously, conditioned on human videos and historical context.
-Inspired by  [27], Human2Robot explores a Spatial UNet and a Spatial-Temporal UNet, working collaboratively to learn from humans.
-In particular, the Spatial UNet extracts features from the robot arm, which is further fed into the Spatial-Temporal UNet for temporal modeling.
-Human2Robot also contains two behavior extractors that estimate position and motion clues using human videos.
-To achieve action prediction, we use the intermediate features from the Spatial-Temporal UNet to predict the action trajectory via a Policy Head.
+In the first stage, we train a Video Prediction Model (VPM) to generate robotic arm videos conditioned on human videos.
+In the second stage, we freeze the VPM and train an action decoder to predict robot actions based on the motion features generated by the VPM.
+In this section, we describe the two-stage training process of our Human2Robot.
+First, we explore training a VPM to generate robot videos based on human videos, implicitly learning the corresponding robotic arm actions from human movements.
+Next, we design an action decoder that aggregates the predictive action representations within the VPM and outputs the corresponding robot actions.
+4.1 Stage1 : Video Prediction Model (VPM)
+We first explored the possibility of generating robot videos from human videos.
+Inspired by  (Hu 2024), our VPM explores a Spatial UNet (S-UNet in short) and a Spatial-Temporal UNet (ST-UNet in short), working collaboratively to learn from humans.
+In particular, the S-UNet extracts features from the robot arm, which is further fed into the ST-UNet for temporal modeling.
+VPM also contains a behavior extractors that estimate position and motion clues from human videos.
 Figure 3 gives an overview of our pipeline.
-Below, we introduce the architecture in detail.
+Below, we introduce the architecture of VPM in detail.
+[…]
 Spatial UNet.
-The Spatial UNet, consisting of 4 upsampling blocks and 4 downsampling blocks, extracts useful clues from the robot arm.
-In particular, each block, known as the Spatial Layer, leverages self-attention to learn features tailored for robot arms, as well as cross-attention with pre-extracted CLIP embedding to incorporate semantic clues.
-The weights of the Spatial-UNet are initialized from those of the Stable Diffusion [48] to ease the learning process.
-The features derived by Spatial-UNet provide a condition for predicting future frames as well as actions.
-While Spatial UNet introduces a similar number of parameters as the denoising UNet, it only extracts features once during the entire process, unlike diffusion-based video generation where each frame undergoes denoising multiple times.
+The S-UNet, consisting of 4 upsampling blocks and 4 downsampling blocks, extracts useful clues from the robotic arm.
+In particular, each block, known as the Spatial Layer (S-Layer in short), leverages self-attention to learn features tailored for robot arms, as well as cross-attention with pre-extracted CLIP embedding to incorporate semantic clues.
+The weights of the S-UNet are initialized from those of the Stable Diffusion (Rombach et al.
+2022) to ease the learning process.
+The features derived by S-UNet provide a condition for predicting future frames.
+While S-UNet introduces a similar number of parameters as the denoising UNet, it only extracts features once during the entire process, unlike diffusion-based video generation where each frame undergoes denoising multiple times.
 Therefore, it does not significantly increase computational overhead during inference.
-Behavior Extractor.
-Recall that we aim to learn from human demonstrations, and thus it is important to extract motion and position information from human videos.
-This is achieved by Behavior Extractor, which contains four convolution layers (4$\times$ kernels, with a stride of 2).
-To focus more on the motion trajectories, we employ two independent Behavior Extractors.
-Behavior Extractor 1 takes the image of the hand corresponding to the robot at the first time step, and Behavior Extractor 2 takes the future image or the entire video of the hand as inputs depending on the training stage, as will be introduced in Section 4.2.
 Spatial-Temporal UNet.
-The Spatial-Temporal UNet aims to predict future frames as well as actions to take simultaneously, requiring the model to learn temporal dynamics.
-Thus, each block of the Spatial-Temporal UNet (ST-UNet in short) has a Spatial Layer which is the same as the one in Spatial UNet (S-UNet in short), followed by an additional Temporal Layer.
+The ST-UNet aims to predict future frames, requiring the model to learn temporal dynamics.
+Thus, each block of the ST-UNet has a S-Layer which is the same as the one in S-UNet, followed by an additional Temporal Layer (T-Layer in short).
 More specifically, the ST-UNet takes the features from the behavior extractors and noise latent as inputs.
-Each Spatial Layer in ST-UNet also concatenates the features of the corresponding layer of the S-UNet, exploring the robot arm as references for prediction.
-The Temporal Layer focuses on temporal modeling so as to produce high-fidelity future frames as well as accurate action trajectories.
-Policy Head.
-We aim to generate future frames as well as actions to take within a single unified framework.
-Drawing from the insights of [62], we propose to divide the UNet mapping function into two distinct components: $F=F_{tail}\cdot F_{head}$.
-We treat $F_{head}$ as a feature extractor for motion information.
-This is achieved by passing the intermediate features from UNet and proprioception data $s_{t}$ to a light-weighted Multilayer Perceptron to predict the actions of the robotic arm.
-Pretrained Encoder and Decoder.
-Human2Robot builds upon Stable Diffusion [48], which consists of an encoder and a decoder.
+In addition, each S-Layer in ST-UNet also concatenates the features of the corresponding layer of the S-UNet, exploring the robot arm as references for prediction.
+The T-Layer focuses on temporal modeling so as to produce high-fidelity future frames.
+VAE Encoder and Decoder.
+Human2Robot builds upon Stable Diffusion (Rombach et al.
+2022), which consists of an encoder and a decoder.
 The encoder turns images into latent embeddings for fast denoising, and the decoder maps the latents back to images.
 During training, both the encoder and decoder are kept frozen.
-4.2 Training
-Our approach consists of two training stages.
-In the first stage, we focus on image generation by taking the first frame of $o^{r}_{t}$, the first frame of human $o^{h}_{t}$ and future frame of human $o^{h}_{t+i}$ as inputs to predict the future frame of robot $o^{r}_{t+i}$.
+Training strategy.
+To generate higher-quality videos, the VPM training process consists of two stages.
+In the first stage, we focus on image generation to acquire basic generative capabilities.
+In the second stage, we train it to generate videos.
+In the first stage, by taking the first frame of robot $o^{r}_{0}$, the first frame of human $o^{h}_{0}$ and future frame of human $o^{h}_{i}\ (i\in[0,T])$ as inputs to predict the future frame of robot $o^{r}_{i}$, where $T$ refer to the maximum length of video generation.
 In this stage, we train the S-UNet, the Behavior Extractors, and the ST-UNet without the temporal layers.
-The spatial layers in both the ST-UNet and S-UNet are instantiated using the powerful open-source Stable Diffusion model (SD)  [48], while the Behavior Extractors are initialized with Gaussian weights, except for the final projection layer, which utilizes zero initialized convolution.
-In the second stage, we concentrate the training effort on the temporal layer.
-In this stage, we take a 30-frame segment of human video $o^{h}_{t:t+n}$ and the first frame $o^{r}_{t}$ of the robot video $o^{r}_{t:t+n}$ as inputs to predict the future robot observation $o^{r}_{t:t+n}$ and action $a_{t:t+n}$.
-The first frame of the robot video is fed into Behavior Extractor 1, while the entire video segment is passed into Behavior Extractor 2, yielding features $x_{1}^{t}$ and $x_{2}^{t:t+n}$, respectively.
-Then, we replicate $x_{1}^{t}$ by n times and add it to $x_{2}^{t:t+n}$, which is subsequently added to noise latent and fed to our ST-UNet.
+The S-Layer in both the ST-UNet and S-UNet are instantiated using the powerful open-source Stable Diffusion model (SD)  (Rombach et al.
+2022), while the Behavior Extractors are initialized with Gaussian weights, except for the final projection layer, which utilizes zero initialized convolution.
+In the second stage, we concentrate the training effort on the temporal layer to generate video.
+In this stage, we take a 30-frame segment of human video $o^{h}_{0:T}$ and the first frame of robot $o^{r}_{0}$ as inputs to predict the future robot observation $o^{r}_{0:T}$.
+The entire human video segment is first passed into Behavior Extractor, which is subsequently added to noise latent and fed to our ST-UNet.
 The weights of the VAE Encoder and Decoder, as well as the CLIP image encoder are frozen all the time.
-During our training process, we find that dense sampling (30fps) results in jittering in generated videos.
-Therefore, we reduce the frame rate to one-fourth of the original, namely cutting the number of frames by three-quarters.
-Optimization Objectives.
-We train Human2Robot with both generation and action prediction objectives, aiming for the model to achieve high-quality video generation and accurate action prediction simultaneously.
-The goal of generation for the model is to predict the applied noise, so the objective of the optimization process is:
-$L_{G}=\mathop{\mathbb{E}_{z_{ts},c,\epsilon,ts}}(||\epsilon-\epsilon_{\theta}(z_{ts},c,ts)||_{2}^{2}),$ (2)
-the parameters of the formula represent the latent state $z_{ts}$ at time step $ts$, which is obtained from Spatial-Temporal UNet.
-The conditioning variable $c$ includes the first frame $o^{r}_{t}$ of the robot and a human video $o^{h}_{t:t+n}$.
+The optimization objectives for this stage are as follows:
+[…]
+the parameters of the formula represent the latent state $z_{ts}$ at time step $ts$, which is obtained from ST-UNet.
+The conditioning variable $c$ includes the first frame $o^{r}_{0}$ of the robot and a human video $o^{h}_{0:T}$.
 $\epsilon$ denotes the added noise, and $\epsilon_{\theta}(z_{ts},c,ts)$ represents the model’s predicted noise.
-Our action is divided into three components: $a_{t}^{pos}\in\mathbb{R}^{3}$ for 3D position, $a_{t}^{rot}\in\mathbb{R}^{6}$ for rotation and $a_{t}^{grip}\in\{0,1\}$ for gripper state.
-We use 6D rotation representation proposed in [73, 31] to guarantee continue representation.
-We use MSE loss $L_{p}$ and $L_{r}$ for position and rotation respectively, and use BCE loss $L_{g}$ for gripper state.
-To balance the generation and action prediction losses, we used a set of balancing weights to control the contribution of each loss to the overall objective.
-$L=L_{G}+\lambda_{p}L_{p}+\lambda_{r}L_{r}+\lambda_{g}L_{g},$ (3)
-where, $\lambda_{p}$, $\lambda_{r}$ and $\lambda_{g}$ are the weights of $L_{p}$, $L_{r}$ and $L_{g}$ respectively.
-4.3 Inference
-With the aforementioned designs, Human2Robot is able to generate high-quality robot videos and perform actions, using observation captured by the camera of the robot and the provided hand video.
-We further introduce a set of strategies that facilitate more effective task execution by the robot arm, as will be introduced below.
-KNN + Human2Robot.
-To avoid the need to explicitly provide human demonstration videos for every prediction, we use a k-nearest neighbors (KNN) approach to identify the most probable task for the current scene and retrieve the human demonstration video corresponding to the closest matching features as the conditioning input to guide the task execution.
-Specifically, we use DINOv2 [42] and CLIP [44] as feature extractors to capture features from the first frame of each robotic arm video in the training set.
-During prediction, we select the $k$ closest features based on the current environment, and the episode with the most frequent and closest match is chosen as the conditioning input, which is depicted in the KNN-Based Inference section of Figure 3.
-Using KNN significantly simplifies the process for known tasks.
-However, for tasks involving different objectives within the same setting—such as pick cube task in Figure 4 and Table 2, as well as for unseen tasks such as handwriting, explicit human video guidance is required to accomplish these tasks effectively.
-Temporal Addition of Actions.
-To mitigate trajectory jittering during sequential prediction, inspired by the approach used in ACT [72], we derive the action not from a single prediction but from a weighted average of multiple prediction steps.
-In particular, we predict a series of actions $a_{t:t+n}$ at time t, which are then combined with the previously predicted actions to compute the aggregated actions $A_{t:t+n}^{t}$.
-This represents the set of actions predicted at time $t$, to be executed from time $t$ to $t+n$, and can be computed as follows:
-$A_{t:t-m+n}^{t}=(1-\gamma)A_{t:t-m+n}^{t-m}+\gamma a_{t:t-m+n},$ (4)
-$A_{t-m+n:t+n}^{t}=a_{t-m+n:t+n},$ (5)
-where $\gamma$ is a hyperparameter, with a larger $\gamma$ indicating a stronger emphasis on the current prediction, and $m$ ($m\leq n$) refers to the length of the action executed after each prediction.
-Specifically, when $m=n$, the inference degenerates into sequential prediction.
-Inference Acceleration.
-During inference, the diffusion model typically goes through $T$ denoising steps to generate the output.
-We explore the effectiveness of using features from different denoising steps in Section 5.4, and find that features obtained from the first half of the denoising stages lead to better results.
-Therefore, to balance speed and quality, we halt the denoising process after obtaining features from the $\frac{T}{4}$ step, directly predict and execute the action, without waiting for the entire denoising process to complete.
-It is worth noting that if we aim to achieve the fastest possible speed, we can directly use the one-step denoising features for prediction, which allows us to predict the next 30 steps of actions within 0.6 seconds.
-Figure 4: Qualitative results. Visualization of generated result and robot action result.
+4.2 Stage2 : Action Decoding.
+After the first stage of training, our VPM model is able to visually predict the corresponding robotic arm actions based on human movements.
+However, fully denoising an entire video is time-consuming, with most of the time spent on reconstructing image-level details, which are unrelated to manipulation.
+Increasingly, recent research (Hu et al.
+2024; Wen et al.
+2024; Zhu et al.
+2025) has shown that the features of a generative model after a single denoising step already contain sufficient motion features to guide the action head’s planning.
+Inspired by these works, we treat the pretrained VPM as a vision encoder to extract the predicted robotic arm action information.
+This means adding intensity-invariant noise (t = K) to the human video and using the first denoised result as the feature.
+As shown in the Figure 5, we can see that the one-step denoising already contains a lot of action and position information.
+Additionally, previous work (Weng et al.
+2024; Xiang et al.
+2023) has emphasized that the upsampling layers of diffusion models often contain more effective information.
+Therefore, we use the outputs of the VPM’s upsampling layers as features; more precisely, we use those from the first upsampling layer.
+In summary, during this stage, we freeze the parameters of the VPM and treat it as a visual encoder to train the subsequent action decoder.
+We use the first upsampling layer output $F_{\text{VPM}}$, after the initial denoising step, as the action prior features for the action decoder.
+Action Decoder.
+Inspired by recent video conditioned works (Hu et al.
+2024; Luo and Lu 2025), our action decoder consists of two parts: the Video Former (Blattmann et al.
+2023b) and the Diffusion Policy (Chi et al.
+2023).
+The Video Former uses a learnable token $Q$ to aggregate the video features $F_{\text{VPM}}$ into a feature $F_{\text{VF}}\in\mathbb{R}^{n\times L}$ with a fixed-length n and specified channel dimension L.
+Formally, this branch can be expressed as follows:
+$F_{\text{VF}}=\text{FFN}(\text{Temp-Attn}(\text{Spat-Attn}(Q,F_{\text{VPM}}))).$ (2)
+We incorporate the aggregated representation $F_{\text{VF}}$ into the diffusion transformer blocks through cross-attention layers.
+The diffusion policy seeks to reconstruct the original actions $a_{0}$ from the noisy actions $a_{k}=\sqrt{\overline{\beta}_{k}}a_{0}+\sqrt{1-\overline{\beta}_{k}}\epsilon$, where $\epsilon$ represents white noise, and $\overline{\beta_{k}}$ is the noisy coefficient at step $k$.
+The optimization objectives for diffusion policy are as follows:
+[…]
+4.3 KNN + Human2Robot.
+To avoid the need to explicitly provide human demonstration videos for seen task, we use a k-nearest neighbors (KNN) approach to identify the most probable task for the current scene.
+We retrieve the human demonstration video corresponding to the closest matching features as the conditioning input to guide the task execution.
+Specifically, we use DINOv2 (Oquab et al.
+2023) and CLIP (Radford et al.
+2021) as feature extractors to capture features from the first frame of each robotic arm video in the training set.
+During prediction, we select the $n$ closest features based on the current environment, and the episode with the most frequent and closest match is chosen as the conditioning input, which is depicted in the Inference section of Figure 3.
 5 Experiments
+Figure 4: Task overview. We train the models on seen tasks and test them on different generalization ability level.
+Push & Pull Pick & Place Rotation Average
+DP (Chi et al. 2023) 50 20 15 28
+XSkill (Xu et al. 2023b) 70 40 50 53
+VPP (Hu et al. 2024) 95 70 75 80
+Action Decoder w. Human 50 10 10 23
+Human2Robot w/o. Pretrain 20 10 0 10
+Human2Robot w. KNN 90 75 80 82
+Human2Robot (ours) 100 90 95 95
+Table 1: Multi-task success rate on basic tasks. Each task is evaluated with 20 trails.
+Generalization XSkill VPP H2R(ours)
+Appearance 0 80 100
+Position 20 50 80
+Instance 0 0 70
+Background 0 0 80
+Combination 0 0 50
+Brand-New 0 0 70
+Table 2: Generalization success rate. Each task is evaluated with 20 trails.
 5.1 Experimental Setups
-Task Definition. As visualized in Figure 4, we conduct tests across the following four tasks:
-  • Push the box. Push the box from left to right.
-  • Pick the cube to the designated plate. Pick up the cube and place it on the bottom or top plate depending on the human demonstration.
-  • Pick up and hold the brush upright. Pick up the brush, hold it upright, and get ready to write.
-  • Handwriting. Write the corresponding character depending on the human demonstration.
-The four carefully selected tasks above are designed to test the capabilities in multiple aspects: (1) The push box task is to test the accuracy in predicting height, position, and trajectory.
-(2) The pick cube task requires the model to understand the task objective in the video.
-(3) The upright task is a challenging fully 6-DOF task.
-(4) The handwriting task is to test the generalization ability.
-Notably, in the handwriting task, the models are evaluated on unseen characters.
-Since there are countless characters in the world, it is impossible to demonstrate all of them.
-Therefore, having such a generalization ability is crucial.
-Thus, this task demands exceptional generalization capabilities, which we have not observed in other models.
-Baselines. We compare with the following baselines:
-  • RVCD Policy: We design a robot-video-conditioned policy $\pi(a_{t}|s_{t},v)$, where $v$ refers to the robot video that indicates the task.
-Specifically, we modify ACT [72] to a video conditioned policy.
-Details can be found in Supplementary materials D.2.
-  • RVCD Policy w.
-Generated video: The framework and weights of this model is the same as RVCD Policy, but the robot video, generated by Human2Robot using human demonstration, is provided during inference.
-  • XSkill [69]: XSkill is a human-video-conditioned policy tries to bridge the embodiment gap between human and robot. Details can be found in Supplementary materials D.2.
-  • Human2Robot w. KNN Human2Robot with our KNN method proposed in Section 4.3, which enables to perform tasks without explicit demonstrations.
-The baselines and Human2Robot utilize the same H&R dataset for training across all tasks except handwriting.
-For the handwriting task, the models are trained on handwriting play data [60] (randomly moving data).
-Tasks Push Box Pick and Place Upright Handwriting
-Apperence Position Instance Background Bg and Task
-Generalization Type Brown Blue Pos1 Pos2 Pingpong Banana Bg1 Bg2 H R
-XSkill [69]
-Human2Robot
-Table 1: Generalization evaluation.
-We evaluate generalization across four different tasks in unseen settings, including variations in appearances, positions, instances, backgrounds, and task targets.
-The scenes for each task are shown in the figure above.
-Detailed process can be found in the supplementary materials C.
-Each setting is evaluated with one trial.
+Task Definition.
+As mentioned in section 3.1, even with our adjustments, current teleoperation systems still cannot collect data for difficult tasks, such as screwing, when paired human and robot arm videos are required.
+Therefore, we mainly focus training and testing on pick-and-place tasks.
+While individual tasks are relatively simple, we aim to demonstrate the generalization ability of our model after training on such tasks.
+This goal aligns with the spirit of learning from humans, that is, to learn from simple tasks and then generalize to more complex or even new tasks.
+In addition to testing on tasks from the training set, we also test on many unseen tasks, including variations in appearance, position, instance, background, task combinations, and entirely new tasks.
+The task scenarios and their corresponding descriptions are shown in the Figure 4.
+Baselines. Since Human2Robot is characterized by learning from humans and video generation pretraining, we compare it with the following baselines:
+  • Diffusion Policy (Chi et al. 2023): A action learning policy using action diffusers with CLIP language conditions. For simplicity, we refer to it as DP.
+  • XSkill (Xu et al. 2023b): A human-video-conditioned policy through self-supervised learning.
+  • Video Prediction Policy (Hu et al. 2024): A language-conditioned policy that uses video prediction for pretraining. For simplicity, we refer to it as VPP.
+  • Action Decoder w.
+Human: Condition the action decoder of Human2Robot on human videos instead of the features extracted by VPM.
+The human videos are processed by ResNet18 (He et al.
+2016) and used as input to action decoder.
+  • Human2Robot w/o. Pretrain: Human2Robot without the video generation pretraining in Section 4.1.
+  • Human2Robot w. KNN: Human2Robot with our KNN method proposed in Section 4.3, which enables to perform tasks without explicit demonstrations.
+Human2Robot Training Details.
+As mentioned in Section 4, we used a two-stage training method.
+In the first stage, we trained a video prediction model, focusing on generating robotic arm videos based on human videos.
+During this stage, we pre-trained on 2,600 task videos, including some long tasks, such as picking up two blocks.
+Training for the first stage took 3 days using 4 NVIDIA A100 GPUs.
+In the second stage, we trained Human2Robot on simple tasks, as the seen tasks shown in the Figure 4.
+This took about 6 hours using 8 NVIDIA A100 GPUs.
+However, for the writing task mentioned in Section 5.4, we train Human2Robot solely on play data of writing for about 6 hours using 8 NVIDIA A100 GPUs.
+It is worth noting that the play data of writing here consists of random movements, without actually writing any characters.
 5.2 Main results
-Push Box Pick Cube Upright Handwriting
-RVCD w. Gen 60 30 35 0
-RVCD 75 45 40 0
-XSkill [69] 70 40 50 0
-Human2Robot w. KNN 85 40 60 0
-Human2Robot 95 85 80 90
-Table 2: Main results. Each task is evaluated with 20 trails.
-Effectiveness.
-As shown in Table 2, Human2Robot achieves a high success rate across all tasks.
-In the pick cube to the designated plate task, half of the test cases require placing the cube on the bottom plate, while the other half require placing it on the top plate.
-Therefore, a success rate above 50% indicates that the model can correctly follow human demonstrations to perform distinct tasks.
-However, all baselines perform poorly on this task, consistently placing the cube on the bottom plate regardless of the given demonstration.
-This proves that only Human2Robot can truly understand human demonstration and execute the correct tasks.
-In addition, all baselines can successfully complete the handwriting task that requires strong generalization exhibitting they have limited generalization capabilities.
-These results highlight the satisfactory performance and strong generalization capabilities of Human2Robot.
-We also test our KNN method proposed in Section 4.3, in a setting where no demonstrations are given.
-The performance of Human2Robot w.
-KNN is shown in Table 2.
-Human2Robot with KNN outperforms all other baselines across all tasks, demonstrating that even without direct demonstrations, Human2Robot can still achieve strong performance.
-However, the KNN method experiences a 10-20% drop in success rate on the push box and upright tasks, which is within an acceptable range.
-Meanwhile, a significant performance drop is observed in the pick cube to the designated plate task.
-This is expected, as the KNN method struggles to select the desired demonstration video when faced with the same initial scene.
-Finally, the KNN method cannot select demonstrations that were not present during training, as these demonstrations are only used during inference.
-As a result, it fails to accomplish the handwriting task.
-As a whole, such competitive results demonstrate the high efficiency and accuracy of Human2Robot in action prediction with episodes retrieved by KNN.
-Quality of Generated Videos.
-Besides predicting actions, Human2Robot can visually simulate future dynamics of robot movements, requiring precise predictions of the motions of the robot arm and its interactions with objects.
-The visual quality of the generated video of Human2Robot is shown in Figure 4.
-However, to evaluate the quality of generated videos, the traditional methods of using video distribution scores presents the following challenges: (1) Since the test set is small, which is common in the robotics domain, typically containing only 5 to 20 test samples, the scores may be unreliable.
-(2) No existing generation methods are directly applicable to our task.
-To this end, we introduce a new visual quality evaluation pipeline: training a RVCD policy as an upper bound and replacing robot-conditioned videos with generated videos during inference.
-The performance drop serves as a proxy for evaluating visual quality.
-As shown in Table 2, the performance of RVCD w.
-Gen achieve a similar score, demonstrating the realism of our video prediction.
+Quantitative Results.
+The comparison on the basic tasks are show in Table 2.
+However, the DP baseline appears to converge primarily on push and pull tasks and performs poorly on the other tasks.
+Although XSkill conditions on human videos, it merely treats them as task labels.
+As a result, it can complete seen tasks but does not fully exploit the information in the human videos, leading to unstable performance and success rates that are 30–50% lower than Human2Robot.
+VPP also employs video‑generation pretraining and attains success rates close to those of Human2Robot.
+Nevertheless, because VPP is language‑conditioned—whereas Human2Robot is video‑conditioned, providing richer, fine‑grained motion cues—Human2Robot still leads by 10–20 percentage points in success rate.
+Overall, our proposed Human2Robot achieves the highest success rate across all tasks, highlighting the benefits of conditioning on human videos coupled with video‑generation pretraining.
+KNN Results.
+As shown in Table 2, Human2Robot with KNN outperforms all other baselines across all tasks, demonstrating that even without direct demonstrations, Human2Robot can still achieve strong performance.
+However, compared to Human2Robot, the KNN method shows a 10–20% decrease in success rate, which we consider to be within an acceptable range.
+Overall, these competitive results demonstrate that Human2Robot is both efficient and accurate on seen tasks.
+Visualizations of Predictive Representations.
+Figure 5: Visualization of VPM results.
+We can observe that that a 1-step denoised result already contains sufficient motion information for downstream tasks.
+In addition, the 30-step (fully denoised) result is very close to the GT Robot video, demonstrating the effective design of our Video Prediction Model (VPM).
+Since we repurpose the video prediction model as a visual encoder and extract predictive representations with a single forward pass, we examine the quality of these representations.
+In Figure 5, we visualize the ground‑truth future alongside 1‑step predictions and 30‑step denoised outputs.
+The visualizations show that a 1-step denoised result already contains sufficient motion information for downstream tasks, validating the soundness of our approach.
+In addition, the 30-step (fully denoised) result is very close to the GT Robot video, demonstrating the effective design of our VPM.
+5.3 Ablation Study
+Effectiveness of VPM.
+Since our H&R dataset contains paired videos of humans and robotic arms, it naturally includes paired human observations and the corresponding robot actions.
+Therefore, we design Action Decoder w.
+Human to predict robot actions directly from human videos.
+Specifically, we use the human video $o^{h}_{0:T}$ as a conditioning input to the action decoder to predict the corresponding robotic actions $a_{0:T}$.
+However, in our experiments, this approach produced highly jittery executions, and the model was insensitive to grasping motions in the human videos, frequently failing to complete grasps.
+As a result, the average success rate was only 23%.
+We attribute this to the complexity of human hand motions: without any prior for learning the correspondence, it is difficult to infer the correct mapping from human videos.
+By contrast, Human2Robot leverages video generation to learn the action correspondence, providing the downstream action decoder with a more reliable motion prior.
+Effectiveness of Video Pretraining.
+We designed Human2Robot w/o.
+Pretraining to assess the contribution of video‑generation pretraining.
+In this variant, we initialize the VPM with the initial parameters described in Section 4.1, freeze the VPM, and train only the downstream action decoder.
+As shown in the table, this approach is almost incapable of completing tasks: it achieves 20% success on the simplest push‑and‑pull tasks and only 10% on pick‑and‑place.
+We attribute this poor performance to Stable Diffusion initialization fails to extract task‑relevant features and, in effect, injects additional noise into the observations—thereby degrading the policy.
+By contrast, the pretrained Human2Robot delivers strong performance, demonstrating the effectiveness of video pretraining.
 5.4 Generalization
+To evaluate generalization ability, we test not only basic generalization types—position, appearance, instance, and background—but also stronger changes: combined and brand-new tasks.
+As shown in Table 2, compared with the baselines, Human2Robot not only achieves substantially higher metrics on position and appearance, but also maintains solid performance in the other four settings where XSkill and VPP fail.
+We attribute this generalization to two factors: (1) unlike the self‑training approach of XSkill, our method leverages paired human–robot data from H&R, enabling direct learning of the correspondence between human hands and robot arms; and (2) unlike VPP, our policy is human-video‑conditioned rather than language‑conditioned, so human videos supply richer, fine‑grained motion cues, allowing strong generalization even when training on simple tasks and environments.
+Appearance Generalization.
+We tested the generalization ability on objects with different colors, materials, and textures.
+Human2Robot and VPP maintained their core capabilities without being affected, achieving success rates of 100% and 80%, respectively.
+However, XSkill could barely complete the tasks, performing only meaningless random movements, which demonstrates its poor generalization ability.
 Position Generalization.
-Human2Robot uses its generative capability to enable the policy to understand human intentions and perform corresponding tasks at different positions.
-As shown in Table 1, Human2Robot achieves better results in different testing positions.
-Appearance generalization.
-Appearance generalization is crucial for the model’s robustness.
-Since Human2Robot is based on Stable Diffusion [48], which contains a vast amount of prior knowledge, it naturally inherits strong generalization capabilities in terms of appearance without any data augmentation.
-As shown in the Table 1, Human2Robot demonstrates excellent generalization across different colors, which are unseen during training.
+Due to small variations in object placement within the dataset, all three models exhibit some degree of positional generalization.
+However, when the positional shift is large, XSkill achieves only a 20% success rate, VPP reaches only 50%, whereas Human2Robot succeeds on 80% of the tested positions.
 Instance Generalization.
-Instance generalization requires the ability to generalize across objects of different sizes and shapes, making it even more critical than appearance generalization.
-As illustrated in the Table 1, thanks to the generative capability of diffusion models, Human2Robot achieves strong performance across various objects.
-Human2Robot successfully completes tasks on unseen instances, demonstrating its powerful generalization ability.
+We tested the generalization ability on different instances.
+During training, all the models only encountered data for picking up blocks and pens.
+We tested whether the models could generalize to picking up objects such as ping pong balls and bananas.
+Due to the different shapes of the instances, their placement positions also varied slightly, adding complexity to the task.
+Human2Robot still achieved a 70% success rate on this task, while the other baselines were unable to complete it.
 Background Generalization.
-Although the training data primarily consists of white backgrounds, as shown in the Figure 4, we find that Human2Robot maintains stable performance even when the background color changes or the scene becomes cluttered.
-As presented in the Table 1, Human2Robot successfully completes tasks across different background, demonstrating its strong capability.
-Task generalization.
-We train the model on play data, which involves moving randomly on the table rather than writing any specific character, and then evaluate its one-shot task execution ability given human handwriting prompts of never-before-seen character.
-To keep the scene clean during data collection, no actual handwriting was performed, resulting in blank backgrounds in the training data, which differ from those in the evaluation phase.
-Consequently, the handwriting task simultaneously challenges the generalization ability of models in both scene adaptation and task execution.
-The satisfying results shown in the Table 1 demonstrate the effectiveness and generalization capability of Human2Robot in task transfer.
-5.4 Discussion
-The Choice of Upsampling Layers for Feature Extraction.
-As shown in Table 3, we examine the impact of extracting features from different UNet upsample layers for action prediction.
-Using the 2nd up-sampling block (Up Index 2) achieves a 90% success rate on seen tasks, while the 1st up-sampling block (Up Index 1) reaches a 95% success rate on unseen tasks.
-The results indicate that the choice of upsampling layer has minimal effect on outcomes.
-Up Index Seen Task Unseen Task
-1 85 95
-2 90 80
-3 80 70
-Table 3: Results of different up-sampling blocks.
-Studying the Choice of Denoising Step.
-As shown in Tab 4, During inference, we fix the upsampling layer to extract features from different denoising steps.
-We find that, due to the MLP being trained with random noise, features from just one denoising step already allow the model to perform well, achieving 80% accuracy on seen tasks and 90% on unseen tasks, with an inference time of 0.6 seconds for 30 step trajectory prediction.
-Features after 7 denoising steps (1/4 of the total steps) yield the highest success rate.
-However, the results are worse at 15 (1/2 total steps) and 22 (3/4 total steps), possibly due to the following reason: the features in the early denoising stages, such as the 1st and 7th denoising stage shown in Figure 5, already capture sufficient spatial information to complete the tasks.
-However, in later denoising stages, the model focuses more on reconstructing detailed visual information, which increases action variance, making the movements jerky of the robot and ultimately reducing the success rate.
-Denoising Step Seen Task Unseen Task Times(s)
-1 80 90 0.6
-7 85 95 3.2
-15 80 80 6.5
-22 70 80 9.3
-Table 4: Results of different denoising steps.
-Figure 5: Visual result of 1 and 7 step DDIM.
-The Effect of Data Scale.
-After achieving strong image generation capabilities after the first training stage using the whole dataset, we train the model using different data scales in the second stage to assess its performance across varying amounts of data.
-As shown in Table 5, we conduct tests across various data scales and discover that even a small amount of data can lead to promising results on previously encountered tasks.
-For instance, using just 5% of the data(116 episodes) is sufficient to achieve a 70% success rate on seen tasks.
-The success rate on unseen tasks increases significantly as the amount of data increases.
-Data Scale Seen Task Unseen Task
-0.05 70 30
-0.1 70 40
-0.2 80 60
-1 85 80
-Table 5: Results of different data scales.
+We also added many irrelevant objects and introduced unseen backgrounds to test the generalization ability.
+We found that both baselines were unable to make correct predictions under these conditions, while Human2Robot was still able to predict correctly, achieving a success rate of 80%.
+Task Combination.
+We believe that a model with strong generalization ability should not only be able to complete tasks it has seen before, but also be capable of handling unseen tasks.
+At the beginning, we think such a model should have the ability to learn from each short task and then be able to complete long tasks composed of these short tasks.
+The challenge in this task lies in the fact that tasks from the previous stage may affect the execution of the subsequent stage.
+For example, we designed a task that involves pulling the plate and placing the cube onto it.
+Since the plate is moved, the placement position of the block changes, introducing significant difficulty.
+Therefore, only Human2Robot, which has access to the corresponding human video, is capable of completing these tasks.
+Brand-New Task.
+Teaching a robot to write is a challenging problem for current learning-based approaches.
+The core issue lies in the vast number of characters, with writing each character being a brand-new task.
+We cannot teach the model to write each character individually, so it must possess strong generalization capabilities.
+To demonstrate the performance of Human2Robot in writing, we created a special dataset where each trajectory involves random movements on a desktop without writing any specific characters.
+We used this data to train the models to learn the correspondence between human and robots, and then during inference, we instructed the models to write brand-new characters, such as ”H” or ”R.” Experimental results show that only Human2Robot can learn the correspondence from meaningless data and successfully complete the task of writing specified character.
 6 Conclusion
-In this paper, we introduced a data collection pipeline for perfectly paired human-robot data and used it to create the H&R dataset.
-Then we presented Human2Robot, a novel end-to-end diffusion framework that simultaneously predicted robot video and actions given a hand video.
-Next, we leveraged a KNN method to perform tasks without explicitly requiring human hand videos as input.
-Human2Robot not only achieved outstanding performance on four carefully selected tasks but also demonstrated its capabilities in five generalization ability tests.
+In this paper, we use VR teleoperation systems to collect perfectly paired human-robot data, and create the H&R dataset.
+We then present Human2Robot, which leverages a video prediction model (VPM) for human-robot alignment, and the robot-dynamics features captured turn out to be effective in guiding action decoding.
+Our evaluations demonstrate that Human2Robot excels in seen tasks and unseen tasks.
 Appendix A System Setup Details
@@
 Toward this goal, we build our H&R system to develop software and establish data collection protocols.
-We build the control system with OpenTeach [28], connecting different hardware components to enable the operator to perform manual tasks and simultaneously monitor the robotic arm’s actions through the VR headset display.
+We build the control system with OpenTeach (Iyer et al.
+2024), connecting different hardware components to enable the operator to perform manual tasks and simultaneously monitor the robotic arm’s actions through the VR headset display.
 Specifically, we set up two similar environments for human operation and robotic arm operation respectively, as shown in Figure 7, and one camera is placed in each operating space.
@@
 Our goal is to ensure that the relative rotations of the human hand and the robotic arm, relative to their initial states, remain consistent.
-To synchronize the rotation of the robotic arm with the human hand, we establish a 3D coordinate system similar to the approach taken by OpenTeach[28].
+To synchronize the rotation of the robotic arm with the human hand, we establish a 3D coordinate system similar to the approach taken by OpenTeach(Iyer et al.
+2024).
 This coordinate system enables us to compute the rotation matrix $M_{h}^{t}$ for the human hand at time $t$.
@@
 where $P$ is the transformation matrix defined from the basis of the coordinate system $M_{r}^{0}$ to that of $M_{h}^{0}$.
-Organize Equation 8,  9, and 10, and the target rotation matrix of the robotic arm can be calculated as follows:
+Organize Equation 6,  7, and 8, and the target rotation matrix of the robotic arm can be calculated as follows:
 $M_{r}^{i}=M_{r}^{0}P^{-1}(M_{h}^{0})^{-1}M_{h}^{i}P,$ (9)
+Appendix C Supplementary Experiment Results
 Figure 8: Overview of the dataset for base tasks.
-B.2 System stability
+C.1 System stability
 Accuracy.
@@
 Figure 9: Visualization of generalization evaluation process. To accelerate inference, we skip the video generation process when performing the handwriting task.
-Appendix C H&R Details
-C.1 Task Details
+Appendix D H&R Details
+D.1 Task Details
 Our H&R dataset contains 8 base tasks and 6 long-horizon tasks. The definition of our tasks in our dataset is listed below.
@@
   • Return to the initial position: Complete the base tasks and return to the initial position.
-C.2 Episode Details
+D.2 Episode Details
 H&R consists of 2,600 episodes, each containing the following information, adhering to the RT-X standards:
@@
   • timestamp.
-Appendix D Experimental Details
-D.1 Experimental Details
-Implementation Details.
-To demonstrate the effectiveness of our approach, we train our model on H&R.
-We used 4 NVIDIA A100 GPUs for training in the first stage and 8 4090 GPUs for the second stage.
-In the first training stage, we use the native image size of 424×240 and conduct training for 15,000 steps and set the total batch size as 128.
-In the second training stage, the temporal layer and policy header are trained for 50,000 steps using 30-frame video sequences and the total batch size is 8.
-The learning rates of the two stages are set to 1e-5, and the weights of each objective are set to $\lambda_{p}=\lambda_{r}=\lambda_{g}=1$.
-During inference, we utilize a DDIM sampler for 30 denoising steps and extract features from the first upsample layer and the 7th denoising step for action prediction.
-The smoothing parameter $\gamma$ is set to 0.5.
-D.2 Baseline Details
-RVCD Policy.
-Here we talk about how we add conditions to ACT [72] in detail.
-We input the observation information of the robotic arm into the encoder.
-Then, we add the features of the next $n$ frames of the human hand as a condition to the position embedding and input them into the decoder.
-Specifically, similar to the method used for processing the original robot images, we input human images into ResNet18 [22] and obtain a feature map of size 300x512.
-We then pass it through a shared MLP to obtain a feature of size 512, which is added to the position embedding and fed into the Transformer decoder.
-Both the ResNet18 [22] and MLP are trainable.
-XSkill.
-We replicated xSkill in a real-world setting, where the authors did not fully open-source the code.
-As a result, we completed the remaining code based on the code from the simulation environment.
-During the skill discovery and labeled dataset phases, we used the default settings, except for adjusting the num frames parameter of the frame sampler from 100 to 50 to reduce GPU memory consumption.
-We train this stage for 300 epochs.
-In the skill transfer composing phase, we found that performance of XSkill was quite poor.
-Although the behavior cloning loss decreased significantly, the diffusion policy barely demonstrated any task execution ability after denoising.
-Since the XSkill [69] paper mentions that any imitation learning policy can be used with the XSkill framework, we replaced the diffusion policy with an action head the same as Human2Robot, a simple MLP, for training.
-During training, we adjusted the prediction horizon to 32, resized the pretraining to 160x120 to maintain consistency with the previous training process, and set the vision feature dimension to 64, in line with the simulation environment.
-We trained this stage for a total of 2000 epochs until the loss showed no significant changes.

### v4

@@
 Human2Robot employs a Video Prediction Model to learn a rich and implicit representation of robot dynamics by generating robot videos from human input, which in turn guides a decoupled action decoder.
-Our real-world experiments demonstrate that this approach not only achieves high performance on seen tasks but also exhibits significant one-shot generalization to novel positions, objects, instances, and even new task categories.
+Our real-world experiments demonstrate that this approach not only achieves high performance on seen tasks but also exhibits one-shot generalization to novel positions, objects, instances, and even new task categories.
 1 Introduction
@@
 To enable robots to assist with diverse real-world problems, it is essential that they develop a similar capability, particularly the ability to learn directly from human demonstrations.
-This has motivated extensive research (Bahl et al.
-2023; Srirama et al.
-2024; Nair et al.
-2022; Wang et al.
-2023; Bahl, Gupta, and Pathak 2022; Smith et al.
-2019) on how to learn from human-provided demonstrations effectively.
+This has motivated extensive research (bahl2023affordances; srirama2024hrp; nair2022rm; wang2023mimicplay; bahl2022human; smith2019avid) on how to learn from human demonstrations effectively.
 Despite recent progress, current approaches face a fundamental generalization gap.
 While they perform well on tasks encountered during training, they often fail entirely when presented with human demonstrations of unseen tasks.
-The prevailing paradigm relies on applying self-supervised (Xu et al.
-2023b) or contrastive learning (Jain et al.
-2024) methods to coarsely aligned human-robot video pairs.
+The prevailing paradigm relies on applying self-supervised (xu2023xskill) or contrastive learning (jain2024vid2robot) methods to coarsely aligned human-robot video pairs.
 However, the lack of fine-grained supervision in existing datasets fundamentally constrains what models can learn.
 Rather than uncovering detailed action-level correspondences, models are often limited to capturing global features or high-level task summaries.
-Consequently, many approaches extract holistic video representations using models such as the Perceiver Resampler (Alayrac et al.
-2022), which compresses entire clips into fixed-length embeddings.
+Consequently, many approaches extract holistic video representations using models such as the Perceiver Resampler (alayrac2022flamingo), which compresses entire clips into fixed-length embeddings.
 This severely limits their ability to model the nuanced frame-level temporal dynamics that are essential for generalization.
@@
 This necessitates a paradigm shift in both learning methodology and data curation.
-In this paper, we propose that the key to unlocking fine-grained human-robot alignment lies in video generation (Ho et al.
-2022; Esser et al.
-2023; Blattmann et al.
-2023a; Wu et al.
-2023; Khachatryan et al.
-2023; Xing et al.
-2024; Tu et al.
-2024a).
+In this paper, we propose that the key to unlocking fine-grained human-robot alignment lies in video generation (vdm; gen1; align; tuneavideo; text2videozero; simda; tu2024motioneditor).
 Instead of merely learning whether two videos are broadly similar or mapping task goals from the human domain to the robot domain, we aim to generate a corresponding robot video directly from a human demonstration.
@@
 Teleoperation.
-Recently, VR-based methodologies (Iyer et al.
-2024; Ding et al.
-2024) have attracted considerable attention due to their cost-effectiveness, efficiency and versatility.
+Recently, VR-based methodologies (iyer2024open; ding2024bunny) have attracted considerable attention due to their cost-effectiveness, efficiency and versatility.
 However, these methods focus on controlling the robot, whereas our goal of using VR is to capture perfectly aligned videos of humans and robots, which are essential for robotic imitation.
 Learning from Human Videos.
-Researchers have recently been attempting to leverage existing human-centric video datasets to enhance robot policy learning (Liu et al.
-2018; Smith et al.
-2020; Chen, Nair, and Finn 2021; Shaw, Bahl, and Pathak 2022; Zeng et al.
-2024).
-Researchers propose to learn representations from human videos to assist in task execution (Xiao et al.
-2022; Wang et al.
-2023; Majumdar et al.
-2023).
-However, these approaches need substantial prior knowledge and struggle to transfer to robots (Nair et al.
-2022; Bahl et al.
-2023; Bahl, Gupta, and Pathak 2022).
-Meanwhile, some human-video-conditioned methods  (Bharadhwaj et al.
-2024; Jain et al.
-2024; Xu et al.
-2023b) have focused on aligning representations across human and robot videos, they are neither efficient nor capable of generalization.
+Researchers have recently been attempting to leverage existing human-centric video datasets to enhance robot policy learning (liu2018imitation; smith2020avid; chen2021learning; videodex; zeng2024learning; zhang2025vlabench).
+Researchers propose to learn representations from human videos to assist in task execution (xiao2022masked; wang2023mimicplay; majumdar2024searchartificialvisualcortex).
+However, these approaches need strong prior knowledge and struggle to transfer to robots (nair2022rm; bahl2023affordances; bahl2022human).
+Meanwhile, some human-video-conditioned methods  (gen2act; jain2024vid2robot; xu2023xskill) have focused on aligning representations across human and robot videos, they are neither efficient nor capable of generalization.
 In contrast, our approach uses paired data and diffusion models to achieve strong generalization capabilities.
 Diffusion Models for Video Generation.
-Current research has achieved remarkable performance of video generation (Ho et al.
-2022; Esser et al.
-2023; Blattmann et al.
-2023a; Wu et al.
-2023; Khachatryan et al.
-2023; Xing et al.
-2024; Tu et al.
-2024a, b, c).
-Moreover, diffusion models for video stylization (Ye et al.
-2025; Liu et al.
-2023) and human image animation (Hu 2024; Tu et al.
-2024a) have also shown its promising potential for visual and motion transfer.
+Current research has achieved remarkable performance of video generation (vdm; gen1; align; tuneavideo; text2videozero; simda; xing2023svformer; xing2025aid; xing2023vidiff; tu2024motioneditor; tu2024motionfollower; tu2024stableanimator).
+Moreover, diffusion models for video stylization (ye2025stylemaster; liu2023stylecrafter) and human image animation (hu2024animate; tu2024motioneditor) have also shown its promising potential for visual and motion transfer.
 Moreover, diffusion model not only performs impressively in generative domains but also shows promising applications in visual understanding task.
-Some notable approaches involve using diffusion models for object detection (Chen et al.
-2023), image segmentation (Xu et al.
-2023a), and visual representation learning (Yu et al.
-2024; Weng et al.
-2024).
+Some notable approaches involve using diffusion models for object detection (diffusiondet; zhang2025diffusionad), image segmentation (xu2023open), and visual representation learning (yu2024representation; weng2024genrec).
 Thus, we aim to utilize the strong capabilities of diffusion model to learn human-robot alignment and leverages the implicitly learned robot-dynamics features to guide the final policy learning.
@@
 We first explored the possibility of generating robot videos from human videos.
-Inspired by  (Hu 2024), our VPM explores a Spatial UNet (S-UNet in short) and a Spatial-Temporal UNet (ST-UNet in short), working collaboratively to learn from humans.
+Inspired by  (hu2024animate), our VPM explores a Spatial UNet (S-UNet in short) and a Spatial-Temporal UNet (ST-UNet in short), working collaboratively to learn from humans.
 In particular, the S-UNet extracts features from the robot arm, which is further fed into the ST-UNet for temporal modeling.
@@
 In particular, each block, known as the Spatial Layer (S-Layer in short), leverages self-attention to learn features tailored for robot arms, as well as cross-attention with pre-extracted CLIP embedding to incorporate semantic clues.
-The weights of the S-UNet are initialized from those of the Stable Diffusion (Rombach et al.
-2022) to ease the learning process.
+The weights of the S-UNet are initialized from those of the Stable Diffusion (rombach2022high) to ease the learning process.
 The features derived by S-UNet provide a condition for predicting future frames.
@@
 VAE Encoder and Decoder.
-Human2Robot builds upon Stable Diffusion (Rombach et al.
-2022), which consists of an encoder and a decoder.
+Human2Robot builds upon Stable Diffusion (rombach2022high), which consists of an encoder and a decoder.
 The encoder turns images into latent embeddings for fast denoising, and the decoder maps the latents back to images.
@@
 In the first stage, we focus on image generation to acquire basic generative capabilities.
-In the second stage, we train it to generate videos.
+Next, we train it to generate videos.
 In the first stage, by taking the first frame of robot $o^{r}_{0}$, the first frame of human $o^{h}_{0}$ and future frame of human $o^{h}_{i}\ (i\in[0,T])$ as inputs to predict the future frame of robot $o^{r}_{i}$, where $T$ refer to the maximum length of video generation.
 In this stage, we train the S-UNet, the Behavior Extractors, and the ST-UNet without the temporal layers.
-The S-Layer in both the ST-UNet and S-UNet are instantiated using the powerful open-source Stable Diffusion model (SD)  (Rombach et al.
-2022), while the Behavior Extractors are initialized with Gaussian weights, except for the final projection layer, which utilizes zero initialized convolution.
+The S-Layer in both the ST-UNet and S-UNet are instantiated using the powerful open-source Stable Diffusion model (SD)  (rombach2022high), while the Behavior Extractors are initialized with Gaussian weights, except for the final projection layer, which utilizes zero initialized convolution.
 In the second stage, we concentrate the training effort on the temporal layer to generate video.
@@
 However, fully denoising an entire video is time-consuming, with most of the time spent on reconstructing image-level details, which are unrelated to manipulation.
-Increasingly, recent research (Hu et al.
-2024; Wen et al.
-2024; Zhu et al.
-2025) has shown that the features of a generative model after a single denoising step already contain sufficient motion features to guide the action head’s planning.
+Increasingly, recent research (hu2024video; wen2024vidman; zhu2025unified) has shown that the features of a generative model after a single denoising step already contain sufficient motion features to guide the action head’s planning.
 Inspired by these works, we treat the pretrained VPM as a vision encoder to extract the predicted robotic arm action information.
@@
 As shown in the Figure 5, we can see that the one-step denoising already contains a lot of action and position information.
-Additionally, previous work (Weng et al.
-2024; Xiang et al.
-2023) has emphasized that the upsampling layers of diffusion models often contain more effective information.
+Additionally, previous work (weng2024genrec; xiang2023denoising) has emphasized that the upsampling layers of diffusion models often contain more effective information.
 Therefore, we use the outputs of the VPM’s upsampling layers as features; more precisely, we use those from the first upsampling layer.
@@
 Action Decoder.
-Inspired by recent video conditioned works (Hu et al.
-2024; Luo and Lu 2025), our action decoder consists of two parts: the Video Former (Blattmann et al.
-2023b) and the Diffusion Policy (Chi et al.
-2023).
+Inspired by recent video conditioned works (hu2024video; luo2025learning), our action decoder consists of two parts: the Video Former (blattmann2023align) and the Diffusion Policy (chi2023diffusion).
 The Video Former uses a learnable token $Q$ to aggregate the video features $F_{\text{VPM}}$ into a feature $F_{\text{VF}}\in\mathbb{R}^{n\times L}$ with a fixed-length n and specified channel dimension L.
@@
 We retrieve the human demonstration video corresponding to the closest matching features as the conditioning input to guide the task execution.
-Specifically, we use DINOv2 (Oquab et al.
-2023) and CLIP (Radford et al.
-2021) as feature extractors to capture features from the first frame of each robotic arm video in the training set.
+Specifically, we use DINOv2 (oquab2023dinov2) and CLIP (radford2021learning) as feature extractors to capture features from the first frame of each robotic arm video in the training set.
 During prediction, we select the $n$ closest features based on the current environment, and the episode with the most frequent and closest match is chosen as the conditioning input, which is depicted in the Inference section of Figure 3.
@@
 Push & Pull Pick & Place Rotation Average
-DP (Chi et al. 2023) 50 20 15 28
-XSkill (Xu et al. 2023b) 70 40 50 53
-VPP (Hu et al. 2024) 95 70 75 80
+DP (chi2023diffusion) 50 20 15 28
+XSkill (xu2023xskill) 70 40 50 53
+VPP (hu2024video) 95 70 75 80
 Action Decoder w. Human 50 10 10 23
@@
 Baselines. Since Human2Robot is characterized by learning from humans and video generation pretraining, we compare it with the following baselines:
-  • Diffusion Policy (Chi et al. 2023): A action learning policy using action diffusers with CLIP language conditions. For simplicity, we refer to it as DP.
-  • XSkill (Xu et al. 2023b): A human-video-conditioned policy through self-supervised learning.
-  • Video Prediction Policy (Hu et al. 2024): A language-conditioned policy that uses video prediction for pretraining. For simplicity, we refer to it as VPP.
+  • Diffusion Policy (chi2023diffusion): A action learning policy using action diffusers with CLIP language conditions. For simplicity, we refer to it as DP.
+  • XSkill (xu2023xskill): A human-video-conditioned policy through self-supervised learning.
+  • Video Prediction Policy (hu2024video): A language-conditioned policy that uses video prediction for pretraining. For simplicity, we refer to it as VPP.
   • Action Decoder w.
 Human: Condition the action decoder of Human2Robot on human videos instead of the features extracted by VPM.
-The human videos are processed by ResNet18 (He et al.
-2016) and used as input to action decoder.
+The human videos are processed by ResNet18 (he2016deep) and used as input to action decoder.
   • Human2Robot w/o. Pretrain: Human2Robot without the video generation pretraining in Section 4.1.
@@
 As a result, the average success rate was only 23%.
-We attribute this to the complexity of human hand motions: without any prior for learning the correspondence, it is difficult to infer the correct mapping from human videos.
+We attribute this to the complexity of human motions: without any prior for learning the correspondence, it is difficult to infer the correct mapping from human videos.
 By contrast, Human2Robot leverages video generation to learn the action correspondence, providing the downstream action decoder with a more reliable motion prior.
@@
 Human2Robot and VPP maintained their core capabilities without being affected, achieving success rates of 100% and 80%, respectively.
-However, XSkill could barely complete the tasks, performing only meaningless random movements, which demonstrates its poor generalization ability.
+However, XSkill could barely complete the tasks, showing limited generalization.
 Position Generalization.
@@
 To demonstrate the performance of Human2Robot in writing, we created a special dataset where each trajectory involves random movements on a desktop without writing any specific characters.
-We used this data to train the models to learn the correspondence between human and robots, and then during inference, we instructed the models to write brand-new characters, such as ”H” or ”R.” Experimental results show that only Human2Robot can learn the correspondence from meaningless data and successfully complete the task of writing specified character.
+We used this data to train the models to learn the correspondence between human and robots, and then during inference, we instructed the models to write brand-new characters, such as ”H” or ”R.” Experimental results show that only Human2Robot can learn the correspondence from meaningless data and complete the task of writing character.
 6 Conclusion
@@
 Our evaluations demonstrate that Human2Robot excels in seen tasks and unseen tasks.
+Acknowledgments.
+This work was supported by the Science and Technology Commission of Shanghai Municipality(No. 24511103100).
 Appendix A System Setup Details
@@
 Toward this goal, we build our H&R system to develop software and establish data collection protocols.
-We build the control system with OpenTeach (Iyer et al.
-2024), connecting different hardware components to enable the operator to perform manual tasks and simultaneously monitor the robotic arm’s actions through the VR headset display.
+We build the control system with OpenTeach (iyer2024open), connecting different hardware components to enable the operator to perform manual tasks and simultaneously monitor the robotic arm’s actions through the VR headset display.
 Specifically, we set up two similar environments for human operation and robotic arm operation respectively, as shown in Figure 7, and one camera is placed in each operating space.
@@
 Our goal is to ensure that the relative rotations of the human hand and the robotic arm, relative to their initial states, remain consistent.
-To synchronize the rotation of the robotic arm with the human hand, we establish a 3D coordinate system similar to the approach taken by OpenTeach(Iyer et al.
-2024).
+To synchronize the rotation of the robotic arm with the human hand, we establish a 3D coordinate system similar to the approach taken by OpenTeach(iyer2024open).
 This coordinate system enables us to compute the rotation matrix $M_{h}^{t}$ for the human hand at time $t$.